### v3

# Microscopic Theory of Cooper-pair fluctuations in a 2D Electron System with Strong Spin-Orbit Scatterings: Crossover to Phase Separation and Large Magnetoresistance

Abstract
A microscopic theory of Cooper-pair fluctuations (CPFs) in a 2D electron system with strong spin-orbit scatterings is presented to account for the observation, at low temperatures, of large magnetoresistance (MR) above a crossover field to superconductivity in electron-doped SrTiO3/LaAlO3 interfaces. It is found that the conventional (diagrammatic) microscopic theory of superconducting fluctuations yields corrections to the normal-state conductivity too small to account for the pronounced MR. However, an extension of the theory, which introduces a proper definition of the CPFs density, reveals an operative mechanism of this intriguing large MR. A crossover to phase separation, due to Coulomb repulsion, between the rarefying unpaired normal-state electron gas and condensates of CPFs in real-space mesoscopic puddles, developed under increasing field, drive a crossover to a large MR. In the framework of this scenario it is possible to amend the state-of-the-art microscopic theory consistently with the time dependent Ginzburg-Landau functional approach, used previously in an extensive account of the experimentally reported effect.
I Introduction
It was shown recently MZPRB2021 ,MZJPC2023 that Cooper-pair fluctuations (CPFs) in a 2D electron system with strong spin-orbit scatterings can lead at low temperatures to pronounced magnetoresistance (MR) above a crossover field to superconductivity. Employing the time dependent Ginzburg-Landau (TDGL) functional approach the model was applied to the high mobility electron systems formed in the electron-doped interfaces between two insulating perovskite oxides—SrTiO3 and LaAlO3 Ohtomo04 ,Caviglia08 , showing good quantitative agreement with a large body of experimental sheet-resistance data obtained under varying gate voltage Mograbi19 (see also RoutPRL2017 ,ManivNatCommun2015 ).
In the present paper we approach the same problem from the point of view of the conventional (diagrammatic) microscopic theory of fluctuations in superconductors, developed by Larkin and Varlamov (LV) LV05 . Consistently with the state-of-the-art microscopic theory of superconducting fluctuations at very low temperatures GalitLarkinPRB01 , Glatzetal2011 , Lopatinetal05 , Khodas2012 , our model calculations yield corrections to the normal-state conductivity, which are too small to account for the observed pronounced MR effect. However, further analysis of the results of the microscopic theory reveals that the density of CPFs, defined consistently with the TDGL approach, is a key function of field and temperature for understanding the large MR phenomenon. It is then found that, in the zero temperature limit at finite magnetic field, the CPFs density diverges within restricted mesoscopic regions in space so that the developing huge local charging under increasing field should eventually drive, due to Coulomb repulsion, the (unpaired) normal-state electrons out of the dense CPFs regions. Under these circumstances the normal-state electrons, which are not affected by CPFs within the bulk of the condensed puddles, hold mutual charge exchange with CPFs via interphase dynamical equilibrium.
It has been, therefore, concluded that during the extended life-time of CPFs in their mesoscopic enclaves, when the fluctuation paraconductivity diminishes, the density of the normal-state conduction electrons is also suppressed (due to charge transfer to the localized CPFs) and so yielding overall large MR. The crossover to this low temperature state of pronounced MR is, therefore, associated with the tendency of the nearly uniform system of CPFs at zero field to formation, under increasing field, of highly inhomogeneous system of condensed CPFs puddles. Following this scenario it is possible to amend the LV microscopic theory, in a consistent manner with the TDGL functional approach presented in Refs.MZPRB2021 , MZJPC2023 .
The paper is organized as follows: In Sec.II we present the model of a 2D electron system with strong spin-orbit scatterings employed in this paper. The model is then applied in Sec.III to the microscopic (zero field) LV theory at very low temperature. Subsequently, in the same section, the results of the microscopic (diagrammatic) theory for the fluctuation conductivity are compared to those of the corresponding TDGL functional approach, revealing important connections between the two approaches. In Sec.IV we present an extension of the microscopic theory of fluctuation conductivity to finite magnetic field, which is further developed in Sec.V by introducing the concept of CPFs density. The phenomenon of field-induced condensation of CPFs at very low temperatures, revealed in analyzing this concept, and its consequent phenomenon of crossover to localization and phase separation of CPFs, are also discussed in Sec.V.  A detailed discussion of the physical ramifications of the main findings of this paper and concluding remarks appear in SecVI.
Figure 1: Right figure:  Leading-order Feynman diagrams of the current-current correlator contributing to the fluctuation conductivity. Small empty circles are bare current vertices, solid lines are impurity average normal-state Green’s functions, wavy lines are fluctuation propagators, small full circles are bare pairing vertices, dashed areas stand for impurity ladders between two electron lines, dashed lines with central crosses are additional impurity renormalizations. Left figure: External (upper row) and internal (middle row) impurity-scattering renormalization of pairing vertices appearing in the Feynman diagrams on the right and in the fluctuation propagator (lower row on the left).
II The model
The model employed in this paper, following Refs. MZPRB2021 ,MZJPC2023 , which have been motivated by the perovskite oxides electronic interface state investigated in Ref. Mograbi19 , consists of a thin rectangular film of disordered electron system, under a strong magnetic field $H$, applied parallel to the conducting plane. Disorder is due to impurity scatterings, including both nonmagnetic potential scattering and spin-orbit scattering Abrikosov62 , Klemm75 . Superconductivity in this system is governed by the interplay between the Zeeman spin splitting energy,$\mu_{B}H$ and the spin-orbit scattering rate, $1/\tau_{SO}\equiv\varepsilon_{SO}/\hbar$ (see a detailed description in early papers dealing with similar 3D systems Maki66 , FuldeMaki70 , WHH66 ), where the corresponding spin-flip scattering processes effectively suppress the pair-breaking effect of the Zeeman spin-splitting. The nonmagnetic potential scatterings influence the CPFs only through the coherence length, i.e. via the effect of the scattering rate on the electronic diffusion constant (see Ref.Maki66 ).
The spin-orbit scattering matrix employed, following Ref.MZPRB2021 , has been a reasonable model for the strong spin-orbit interaction of the Ti 3d conduction electrons with lattice ions in SrTiO3(CavigliaPRL2010 ). For the (111) LaAlO3/SrTiO3 interface employed in Ref.Mograbi19 the spin-relaxation time $\tau_{SO}$, was found (see Ref.RoutPRL2017 ) to follow the (Elliott-Yafet) relation CavigliaPRL2010 ,ZuticRMP2004 : $\tau_{SO}\sim\tau_{p}$, where $\tau_{p}$ is the elastic momentum relaxation time, in a region of gate voltages where crossovers from superconductivity to large MR have been observed Mograbi19 .
Under these circumstances the use of $\tau_{SO}$ as a single relaxation time in our calculations seems reasonable. However, one can readily show, on the basis of results published long time ago Maki66 , WHH66 (and reproduced for a 2D system in unpublished ), that our model adequately describes also the usual situations where the non-magnetic potential scatterings dominate the relaxation processes, the only modification due to the non-magnetic scatterings appears through the electronic diffusion constant controlling the fluctuation kinetic energy.
[…]
Our model is similar to the models employed in both Refs.Lopatinetal05 and Khodas2012 , but differs in an important aspect. In our model impurity scatterings are dominated by spin-orbit interactions whereas in Refs.Lopatinetal05 , Khodas2012 spin-orbit interaction was absent.
III TDGL functional approach vs. microscopic theory at zero field
III.1 The microscopic Larkin-Varlamov (diagrammatic) theory
In this subsection we introduce the formalism, employed by LV for the zero magnetic field situation at temperatures above the zero-field transition temperature $T_{c0}$, in a general form which allows extension to finite field at low temperatures well below $T_{c0}$. It is implicitly assumed, throughout this paper, that the critical shift parameter $\varepsilon$ (or $\varepsilon_{H}$ for the finite-field situations, see below for more details) includes high-order terms in the Gorkov GL expansion, self-consistently in the interaction between fluctuations UllDor90 ,UllDor91 , as done in Ref.MZPRB2021 .
As indicated in Sec.II, the incorporation of such interactions self-consistently into the equation determining the ”critical” shift parameter avoids it vanishing at any field and temperature and so, in particular, removes the quantum critical point (Lopatinetal05 , Khodas2012 ). Our motivation in selecting this approach is the absence of genuine zero resistance in the experimental data reported in Ref.Mograbi19 , even at zero field.
III.1.1 The Aslamazov-Larkin paraconductivity diagram
The principal contribution to the conductivity due to CPFs (paraconductivity) in the microscopic (diagrammatic) theory is associated with the Aslamazov-Larkin (AL) diagram AL68 (Number 1 in Fig.1, see also Fig.2). The corresponding static Aslamazov-Larkin conductivity $\sigma_{AL}^{LV}$ is obtained from the retarded current-current correlator $Q_{xx}^{LV\left(1\right)}\left(\omega\right)$ : $\sigma_{AL}^{LV}=\lim_{\omega\rightarrow 0}\frac{i}{\omega}\left[Q_{xx}^{LV\left(1\right)}\left(\omega\right)-Q_{xx}^{LV\left(1\right)}\left(0\right)\right]$, where $Q_{xx}^{LV\left(1\right)}\left(\omega\right)$ is the analytic continuation of the Matzubara imaginary time-ordered correlator, i.e.: $Q_{xx}^{LV\left(1\right)}\left(\omega\right)=Q_{xx}^{LV(1)}\left(i\Omega_{\nu}\rightarrow\omega+i\delta\right)$. The corresponding time-ordered correlator in imaginary frequency representation can be written as LV05 :
$Q_{xx}^{LV(1)}\left(i\Omega_{\nu}\right)=-4e^{2}k_{B}T\sum\limits_{k=-\infty}^{\infty}\frac{1}{d}\left(\frac{1}{2\pi}\right)^{2}\int d^{2}q$ (1)
×$\times\mathcal{D}\left(q,\Omega_{k}+\Omega_{\nu}\right)B_{x}\left(q,\Omega_{k},\Omega_{\nu}\right)\mathcal{D}\left(q,\Omega_{k}\right)B_{x}\left(q,\Omega_{k},\Omega_{\nu}\right)$
where:
$\mathcal{D}\left(q,\Omega_{k}\right)=\frac{1}{N_{2D}\left[\varepsilon+\frac{\pi\hbar}{8k_{B}T}\left(Dq^{2}+\left|\Omega_{k}\right|\right)\right]}$ (2)
is the fluctuation propagator in wavenumber-(Matsubara) frequency representation, $\varepsilon=\ln\left(T/T_{c0}\right)$, $T_{c0}$ is the zero-field transition temperature, $D=\tau_{SO}v_{F}^{2}/2$ is the electronic diffusion coefficient, and $v_{F}=\hbar k_{F}/m^{\ast}$ is the Fermi velocity. Here $\Omega_{k}=2kk_{B}T/\hbar,\Omega_{\nu}=2\nu k_{B}T/\hbar,$ $k=0,\pm 1,\pm 2,...,$    $\nu=0,1,2,....$ are bosonic Matsubara frequencies and $N_{2D}=m^{\ast}/2\pi\hbar^{2}$ is the single-electron density of states (DOS), with an effective mass $m^{\ast}$. It should be also noted that Eq.2 is a small-wavenumber approximation of the fluctuation propagator at zero field, derived in Ref.MZPRB2021 (see Eq.41).
In Eq.1 the effective current vertex part (see Fig.2) is given by:
$B_{x}\left(q,\Omega_{k},\Omega_{\nu}\right)$ $=$ $k_{B}T\sum\limits_{n=-\infty}^{\infty}\lambda\left(q,\omega_{n}+\Omega_{\nu},\Omega_{k}-\omega_{n}\right)\lambda\left(q,\omega_{n},\Omega_{k}-\omega_{n}\right)\times$
$\left(\frac{1}{2\pi}\right)^{2}\int d^{2}pG\left(\mathbf{p},\omega_{n}+\Omega_{\nu}\right)G\left(\mathbf{p},\omega_{n}\right)G\left(\mathbf{q-p},\Omega_{k}-\omega_{n}\right)v_{x}\left(\mathbf{p}\right)$
where $\omega_{n}=\left(2n+1\right)k_{B}T/\hbar,n=0,\pm 1,\pm 2,...$ , is a fermionic Matsubara frequency, and $\lambda$ stands for a three-leg vertex of two electron lines and a fluctuation line, renormalized by impurity-scattering ladder (see left Fig.1). A single-electron line corresponds to the Green’s function: $G\left(\mathbf{p},\omega_{n}\right)=1/\left[i\hbar\widetilde{\omega}_{n}-\hbar^{2}\left(p^{2}-k_{F}^{2}\right)/2m^{\ast}\right]$, whereas the ”bare” current vertex is: $v_{x}\left(\mathbf{p}\right)=\hbar p_{x}/m^{\ast}$.
At zero magnetic field, following LV, we have:
$\lambda\left(q,\omega_{n},\Omega_{k}-\omega_{n}\right)=\frac{\left|\widetilde{\omega}_{n}-\left(\widetilde{\Omega_{k}-\omega}_{n}\right)\right|}{\left|2\omega_{n}-\Omega_{k}\right|+\frac{v_{F}^{2}q^{2}}{2\tau_{SO}\left|\widetilde{\omega}_{n}-\left(\widetilde{\Omega_{k}-\omega}_{n}\right)\right|^{2}}}$ (4)
where $\widetilde{\omega}_{n}=\omega_{n}+\left(1/2\tau_{SO}\right)sign\left(\omega_{n}\right)$, whereas $B_{x}\left(q,\Omega_{k};\Omega_{\nu}\right)$ is approximated by taking: $\Omega_{\nu}=0,$ $\Omega_{k}=0$, so that:
$B_{x}\left(q,\Omega_{k},\Omega_{\nu}\right)\simeq B_{x}\left(q,0,0\right)=k_{B}T\sum\limits_{n=-\infty}^{\infty}\lambda^{2}\left(q,\omega_{n},-\omega_{n}\right)$
$\left(\frac{1}{2\pi}\right)^{2}\int d^{2}pG\left(\mathbf{p},\omega_{n}\right)G\left(\mathbf{p},\omega_{n}\right)G\left(\mathbf{q-p},-\omega_{n}\right)v_{x}\left(\mathbf{p}\right)$
and:
$\lambda\left(q,\omega_{n},-\omega_{n}\right)=\frac{\left|\widetilde{\omega}_{n}\right|}{\left|\omega_{n}\right|+\frac{v_{F}^{2}q^{2}}{4\tau_{SO}\left|2\widetilde{\omega}_{n}\right|^{2}}}\simeq\frac{\left|\widetilde{\omega}_{n}\right|}{\left|\omega_{n}\right|+\frac{1}{2}Dq^{2}}$ (5)
Performing the integration over the electronic wavevector and the Matsubara frequency summation we find:
$B_{x}\left(q,0,0\right)\approx-N_{2D}\tau_{SO}v_{F}^{2}q_{x}\frac{1}{4\pi k_{B}T}\psi^{\prime}\left(\frac{1}{2}+\frac{D}{4\pi k_{B}T}q^{2}\right)$ (6)
where $\psi$ is digamma function.
Neglecting the $q$ dependence, i.e. taking $B_{x}\left(q\rightarrow 0,0,0\right)\rightarrow-2N_{2D}\eta_{\left(2\right)}q_{x}$, with:
$\eta_{\left(2\right)}=\frac{\pi\hbar D}{8k_{B}T}$ (7)
the corresponding expression for the zero-field AL conductivity is found to be identical to the LV integral form of the well-known result:
$\sigma_{AL}^{LV}=\frac{e^{2}}{8\hbar d}\int d\left(\eta_{\left(2\right)}q^{2}\right)\frac{\left(\eta_{\left(2\right)}q^{2}\right)}{\left(\varepsilon+\eta_{\left(2\right)}q^{2}\right)^{3}}$
Figure 2: Lower figure: AL diagram consisting of effective current vertices ($B_{x}$) and fluctuation propagators ($\mathcal{D}$). Upper figure: DOS conductivity diagram illustrating the mechanism in which electron time of coherence is suppressed due to electron scattering by background electrons via virtual exchange of CPF with $q\neq 0$ and $I\neq 0$.
III.1.2 The DOS conductivity diagrams
The next leading contributions to the fluctuation conductivity in the LV scheme corresponds to four bubble-shaped diagrams (numbers 5-8 in Fig.1), termed density-of-states diagrams, due to the fluctuation self-energy insertion to single-electron lines. Other, topologically distinct diagrams of the same order in the fluctuations propagators, (see in Fig.1 the diagrams numbered 2-4), well-known as the Maki-Thompson MakiPTP1968 ,ThompsonPRB1970 , type of diagrams, have been disregarded in our model since the strong spin-orbit scatterings, which characterize the SrTiO3/LaAlO3 interfaces under consideration, are known to suppress their overall contribution LV05 (See Appendix A). Furthermore, as explained in LV, the DOS diagrams 9,10 can also be neglected, however at very low temperature and close to the quantum critical point they dominate the conductivity (see Refs.Lopatinetal05 , Glatzetal2011 , Khodas2012 ). We will return to the issue of quantum critical fluctuations in Sec.VI.
The current-current correlator corresponding, e.g. to diagram 5 (see also Fig.2 upper part) is written as:
$Q_{xx}^{LV(5)}\left(i\Omega_{\nu}\right)=-2e^{2}k_{B}T\sum\limits_{k=-\infty}^{\infty}\frac{1}{d}\left(\frac{1}{2\pi}\right)^{2}$
×$\times\int d^{2}q\Sigma_{xx}^{(5)}\left(q,\Omega_{k},\Omega_{\nu}\right)\mathcal{D}\left(q,\Omega_{k}\right)$ (8)
where:
$\Sigma_{xx}^{(5)}\left(q,\Omega_{k},\Omega_{\nu}\right)=k_{B}T\sum\limits_{n=-\infty}^{\infty}\lambda^{2}\left(q,\omega_{n},\Omega_{k}-\omega_{n}\right)$
×$\times I_{xx}^{(5)}\left(\mathbf{q},\omega_{n},\Omega_{k},\Omega_{\nu}\right)$ (9)
and:
$I_{xx}^{(5)}\left(\mathbf{q},\omega_{n},\Omega_{k},\Omega_{\nu}\right)=\left(\frac{1}{2\pi}\right)^{2}\int d^{2}pG^{2}\left(\mathbf{p},\omega_{n}\right)$
$G\left(\mathbf{q-p},\Omega_{k}-\omega_{n}\right)G\left(\mathbf{p},\omega_{n}+\Omega_{\nu}\right)v_{x}^{2}\left(\mathbf{p}\right)$ (10)
Neglecting dynamical (quantum critical) fluctuations by restricting consideration only to the single term with $\Omega_{k}=0$ in the summation over $k$ in Eq.8, performing the integration over the electron wavevector in Eq.10 and the Matsubara frequency summation in Eq.9, we get after the analytic continuation: $\Omega_{\nu}\rightarrow-i\omega$:
$\Sigma_{xx}^{(5)}\left(q,\Omega_{k}=0,\Omega_{\nu}\rightarrow-i\omega\right)\approx-\frac{i\omega\tau_{SO}^{2}v_{F}^{2}N_{2D}}{8\pi k_{B}T\hbar}\times$
$\left[\psi^{\prime}\left(\frac{1}{2}+\frac{\hbar}{4\pi k_{B}T\tau_{SO}}\right)-\frac{3\hbar}{4\pi k_{B}T\tau_{SO}}\psi^{\prime\prime}\left(\frac{1}{2}+\frac{\hbar Dq^{2}}{4\pi k_{B}T}\right)\right]$
$\approx\frac{i\omega DN_{2D}}{4\pi k_{B}T}\frac{3}{4\pi k_{B}T}\psi^{\prime\prime}\left(\frac{1}{2}+\frac{\hbar Dq^{2}}{4\pi k_{B}T}\right)$ (11)
Further simplification, consistent with the procedure used in evaluating the AL contribution, is achieved by neglecting the $q$ dependence of $\Sigma_{xx}^{(5)}\left(q,\Omega_{k}=0,\Omega_{\nu}\rightarrow-i\omega\right)$ in Eq.8, that is:
$\Sigma_{xx}^{(5)}\left(q\rightarrow 0,\Omega_{k}=0,\Omega_{\nu}\rightarrow-i\omega\right)\rightarrow\frac{i\omega DN_{2D}}{4\pi k_{B}T}\frac{3}{4\pi k_{B}T}\psi^{\prime\prime}\left(\frac{1}{2}\right)$
which yields for the corresponding contribution to the fluctuation conductivity:
$\sigma_{xx}^{LV\left(5\right)}=-\lim_{\omega\rightarrow 0}\frac{Q_{xx}^{LV\left(5\right)}\left(\omega\right)}{i\omega}\simeq-\frac{3e^{2}}{4\hbar d}\frac{14\zeta\left(3\right)}{\pi^{4}}\int\frac{d\left(\eta_{\left(2\right)}q^{2}\right)}{\varepsilon+\left(\eta_{\left(2\right)}q^{2}\right)}$
where we have used the identity: $\psi^{\prime\prime}\left(1/2\right)=-\sum\limits_{n=0}^{\infty}1/\left(n+1/2\right)^{3}=-14\zeta\left(3\right)$.
An identical result is obtained for diagram No.6 (see Fig.1), i.e.: $\sigma_{xx}^{LV\left(6\right)}=\sigma_{xx}^{LV\left(5\right)}$.  A similar method of calculation yields for the 7-th diagram (see also Refs.AltshulerJETP1983 , Dorin-etal-PRB1993 ):
$\sigma_{xx}^{LV\left(7\right)}\simeq\frac{e^{2}}{4\hbar d}\frac{14\zeta\left(3\right)}{\pi^{4}}\int\frac{d\left(\eta_{\left(2\right)}q^{2}\right)}{\varepsilon+\left(\eta_{\left(2\right)}q^{2}\right)}=-\frac{1}{3}\sigma_{xx}^{LV\left(5\right)}$
so that together with the identity: $\sigma_{xx}^{LV\left(7\right)}=\sigma_{xx}^{LV\left(8\right)}$ one finds:
$\sigma_{xx}^{LV\left(5+6+7+8\right)}=2\sigma_{xx}^{LV\left(5+7\right)}=$
$-\frac{e^{2}}{\hbar d}\left(\frac{14\zeta\left(3\right)}{\pi^{4}}\right)\int\frac{d\left(\eta_{\left(2\right)}q^{2}\right)}{\varepsilon+\left(\eta_{\left(2\right)}q^{2}\right)}$ (12)
III.2 Comparison with the TDGL functional approach
III.2.1 The Aslamazov-Larkin paraconductivity
The TDGL functional $\mathfrak{L}\left(\Delta,\mathbf{A}\right)$ of the order parameter $\Delta\left(\boldsymbol{r},t\right)$ and vector potential $\mathbf{A}\left(\boldsymbol{r},t\right)$ determines the Cooper-pairs current density FuldeMaki70 :
$\mathbf{j}\left(\mathbf{r},t\right)=-c\frac{\partial\mathfrak{L}\left(\Delta\left(\mathbf{r},t\right),\mathbf{A}\left(\mathbf{r},t\right)\right)}{\partial\mathbf{A}\left(\mathbf{r},t\right)}$ (13)
responsible for the AL paraconductivity (see Appendix B).
In this approach the entire underlying information about the thin film of pairing electrons system is incorporated in the inverse fluctuation propagator (in wavevector-frequency representation) $D^{-1}\left(\mathbf{q+}2e\mathbf{A/\hbar},-i\Omega\right)$, mediating between the order parameter and the GL functional. In the Gaussian approximation the relation is quadratic, i.e.:
$\mathfrak{L}\left(\Delta,\mathbf{A}\right)$ $=$ $\frac{1}{d}\int\frac{d^{2}q}{\left(2\pi\right)^{2}}\left(\frac{1}{2\pi}\right)\int d\Omega\left|\Delta\left(\mathbf{q},\Omega\right)\right|^{2}\times$ (14)
$\mathcal{D}^{-1}\left(\mathbf{q+}2e\mathbf{A/\hbar},-i\Omega\right)$
so that the coupling to the external electromagnetic field takes place directly through the vertex of the Cooper-pair current, defined in Eq.(13).
The corresponding AL time-ordered current-current correlator is given by:
$Q_{AL}\left(i\Omega_{\nu}\right)=\left(4eN_{2D}D\right)^{2}d^{-1}\left(\frac{1}{2\pi}\right)^{2}\int d^{2}qq_{x}^{2}k_{B}T$ (15)
$\sum\limits_{k=-\infty}^{\infty}C\left(q,\Omega_{k}+\Omega_{\nu}\right)\mathcal{D}\left(q,\Omega_{k}+\Omega_{\nu}\right)C\left(q,\Omega_{k}\right)\mathcal{D}\left(q,\Omega_{k}\right)$
where $\Omega_{k}=2kk_{B}T/\hbar,\Omega_{\nu}=2\nu k_{B}T/\hbar,$ $k=0,\pm 1,\pm 2,...,$    $\nu=0,1,2,....$ are bosonic Matsubara frequencies. Here, like in Sec.IIIA1, the electrical current is generated along the $x$ axis, $q_{z},q_{x}$ are the fluctuation (in-plane) wave-vector components along the magnetic and electric field directions, respectively, and $q^{2}\equiv q_{z}^{2}+q_{x}^{2}$.
The fluctuation propagator $\mathcal{D}\left(q,\Omega_{k}\right)$ and its corresponding effective current vertex $C\left(q,\Omega_{k}\right)$, derived in Ref.MZPRB2021 without restriction to small values of $q$, are given by:
$\mathcal{D}\left(q,\Omega_{k}\right)$ $=$ $\frac{1}{N_{2D}\Phi\left(x+\left|k\right|/2\right)},$
$C\left(q,\Omega_{k}\right)$ $=$ $\frac{\Phi^{\prime}\left(x+\left|k\right|/2\right)}{4\pi k_{B}T}$ (16)
where for zero field:
$\Phi\left(x+\left|k\right|/2\right)=\varepsilon+\psi\left(1/2+x+\left|k\right|/2\right)-\psi\left(1/2\right)$ (17)
and: $x=\hbar Dq^{2}/4\pi k_{B}T$.  Here we note that $\mathcal{D}\left(q,\Omega_{k}\right)$ in Eq.16 reduces to Eq.2 following the linearization of $\Phi\left(x+\left|k\right|/2\right)$.
We now define a normalized effective current vertex analogous to $B_{x}\left(q,0,0\right)$:
$C_{x}\left(q,\Omega_{k}\right)$$\equiv$ $2N_{2D}Dq_{x}C\left(q,\Omega_{k}\right)=2N_{2D}D\frac{\Phi^{\prime}\left(x+\left|k\right|/2\right)}{4\pi k_{B}T}q_{x}$ (18)
$=$ $2N_{2D}D\frac{\psi^{\prime}\left(\frac{1}{2}+\frac{\hbar\left|\Omega_{k}\right|}{4\pi k_{B}T}+\frac{\hbar Dq^{2}}{4\pi k_{B}T}\right)}{4\pi k_{B}T}q_{x}$
with the help of which Eq.15 is rewritten in a form similar to Eq.1, i.e.:
$Q_{AL}\left(i\Omega_{\nu}\right)\simeq 4e^{2}k_{B}T\sum\limits_{k=-\infty}^{\infty}\frac{1}{d}\left(\frac{1}{2\pi}\right)^{2}\times$ (19)
$\int d^{2}q\mathcal{D}\left(q,\Omega_{k}+\Omega_{\nu}\right)C_{x}\left(q,\Omega_{k}\right)\mathcal{D}\left(q,\Omega_{k}\right)C_{x}\left(q,\Omega_{k}\right)$
Taking the static limit, $\Omega_{k}\rightarrow 0$ Eq.18 reduces to:
$C_{x}\left(q,\Omega_{k}\rightarrow 0\right)$$\rightarrow$ $2N_{2D}D\frac{\psi^{\prime}\left(\frac{1}{2}+\frac{\hbar Dq^{2}}{4\pi k_{B}T}\right)}{4\pi k_{B}T}q_{x}$ (20)
$=$ $-B_{x}\left(q,0,0\right)$
The complete agreement between the expressions Eqs.6 and 20, clearly indicates that the effect of the infinite set of ladder diagrams which renormalize the pairing vertices outside the fluctuation propagators (Cooperon insertions) in the AL diagram, within the LV microscopic approach, is fully consistent with the impurity-scatterings effect introduced to the fluctuation current vertex through its relation to the fluctuation propagator within our TDGL functional approach. The corresponding consistency equation between the renormalized current vertex and the fluctuation propagator, inherent to the TDGL functional approach, should also be satisfied within the fully microscopic approach.
Within our TDGL functional approach we note that: $C_{x}\left(q,\Omega_{k}\rightarrow 0\right)=\left(v_{F}^{2}\tau_{SO}/4\pi k_{B}T\right)N_{2D}\Phi^{\prime}\left(x\right)q_{x}$, whereas: $N_{2D}\Phi^{\prime}\left(x\right)=\left(8\pi k_{B}T/\hbar v_{F}^{2}\tau_{SO}\right)\left[\partial/\partial\left(q^{2}\right)\right]\mathcal{D}^{-1}\left(q,0\right)$, so that:
$C_{x}\left(q,0\right)=\frac{2q_{x}}{\hbar}\frac{\partial}{\partial\left(q^{2}\right)}\mathcal{D}^{-1}\left(q,0\right)=-B_{x}\left(q,0,0\right)$ (21)
Equation 21 relates the (Cooperon) external-vertex insertions to the internal-vertex insertions of ladder diagrams introduced in the calculation of the fluctuation propagator. In our TDGL approach it appears naturally, directly from the inverse fluctuation propagator, without taking any additional measure. It reflects the fundamental variation principle, satisfied by the electromagnetically modified GL free energy functional with respect to the vector potential, see Eq.102 in Appendix B.
III.2.2 The DOS conductivity
The basic observable used in evaluating the DOS conductivity within the TDGL approach is the CPFs density:
$n_{s}=\frac{1}{d}\frac{1}{\left(2\pi\right)^{2}}\int\left\langle\left|\phi\left(\mathbf{q}\right)\right|^{2}\right\rangle d^{2}q$ (22)
with the Cooper-pair momentum distribution function $\left\langle\left|\phi\left(\mathbf{q}\right)\right|^{2}\right\rangle$ derived by using the frequency-dependent GL functional, Eq.(14). This is done by rewriting Eq.(14) in terms of the frequency and wavenumber representations GL wavefunctions $\phi\left(\mathbf{q},\Omega\right)$, after analytic continuation to real frequencies $i\Omega_{\mu}\rightarrow\Omega$, i.e.:
$\mathfrak{L}\left(\Delta\right)$ $=$ $\frac{1}{d}\int\frac{d^{2}q}{\left(2\pi\right)^{2}}\int\frac{d\Omega}{2\pi}\left|\Delta\left(\mathbf{q},\Omega\right)\right|^{2}\mathcal{D}^{-1}\left(q,-i\Omega\right)$
$=$ $\frac{1}{d}\int\frac{d^{2}q}{\left(2\pi\right)^{2}}\int\frac{d\Omega}{2\pi}\left|\phi\left(\mathbf{q},\Omega\right)\right|^{2}L^{-1}\left(q,\Omega\right)=\mathfrak{L}\left(\phi\right)$
under the normalization relations:
$\mathcal{D}\left(q,-i\Omega\right)=\left(\frac{\alpha k_{B}T}{N_{2D}}\right)L\left(q,\Omega\right)$ (24)
Here the constant $\alpha$ (i.e. independent of $q$ and $\Omega$), determines the normalization of the wavefunctions from the corresponding components of the order parameter through the reciprocal relations (see Appendix C):
$\left|\phi\left(\mathbf{q},\Omega\right)\right|^{2}=\left(\frac{N_{2D}}{\alpha k_{B}T}\right)\left|\Delta\left(\mathbf{q},\Omega\right)\right|^{2}$
Following LV, the resulting expression for the momentum distribution function is obtained by exploiting the Langevin force technique in the TDGL equation, which leads to:
$\left\langle\left|\phi\left(\mathbf{q}\right)\right|^{2}\right\rangle=2k_{B}T\gamma_{GL}\int\frac{d\left(\hbar\Omega\right)}{2\pi}\left|L\left(q,\Omega\right)\right|^{2}$
where for small $q$ values:
$L\left(q,\Omega\right)^{-1}=k_{B}T\alpha\left(\varepsilon+\xi\left(T\right)^{2}q^{2}\right)-i\gamma_{GL}\hbar\Omega$ (25)
with the CPF coherence length $\xi\left(T\right)$ and dimensionless life-time parameter, $\gamma_{GL}$, are given by:
$\xi\left(T\right)=\sqrt{\frac{\pi\hbar}{8k_{B}T}D},\gamma_{GL}=\frac{\pi\alpha}{8}$ (26)
Note that the actual CPF life-time can be found from the pole of the GL propagator $L\left(q,\Omega\right)$ to be:
$\tau_{GL}\left(q\right)=\frac{\gamma_{GL}}{k_{B}T\alpha\left(\varepsilon+\xi\left(T\right)^{2}q^{2}\right)}$ (27)
Performing the frequency integration we find for the momentum distribution function:
$\left\langle\left|\phi\left(\mathbf{q}\right)\right|^{2}\right\rangle=\frac{1}{\alpha}\frac{1}{\varepsilon+\xi\left(T\right)^{2}q^{2}}$ (28)
At this point one notes that Eq.12 for the LV DOS conductivity has the basic structure of an effective Drude formula, originally proposed by LV, and more recently used within the TDGL approach in Ref.MZJPC2023 , that is:
$\sigma_{xx}^{LV\left(5+6+7+8\right)}\propto-\left(\frac{e^{2}}{d\hbar}\right)\left(\frac{4}{\pi^{2}}\right)\frac{k_{B}T}{E_{F}}\frac{1}{\alpha}\int\frac{d\left(\xi\left(T\right)^{2}q^{2}\right)}{\varepsilon+\xi\left(T\right)^{2}q^{2}}$
$=-2n_{s}\frac{e^{2}}{m^{\ast}}\tau_{SO}\equiv\sigma_{DOS}^{TDGL}$ (29)
where $n_{s}$ is given by Eq.22 and the momentum distribution function by Eq.28. Note also that Eq.12 originates in two equal contributions from two groups of two diagrams shown in Fig.1; diagrams $\left(5,7\right)$, and diagrams $\left(6,8\right)$. Thus, identifying Eq.29 with the basic contribution to the LV DOS conductivity, that is:
$\sigma_{DOS}^{TDGL}=\sigma_{xx}^{LV\left(5+7\right)}=\sigma_{xx}^{LV\left(6+8\right)}$ (30)
the normalization constant should read:
$\alpha=\frac{4\pi^{2}}{7\zeta\left(3\right)}\frac{k_{B}T}{E_{F}}$ (31)
This is exactly the expression for $\alpha$ obtained in the clean limit (i.e. for $\xi\left(T\right)=\sqrt{7\zeta\left(3\right)/8}\left(\hbar v_{F}/2\pi k_{B}T\right)$) by requiring the GL propagator in Eq.25 to have the Schrodinger-like form with the Cooper-pair mass equals twice the free electron mass (see Appendix C). In the dirty-limit situation under study here one therefore evaluates the momentum distribution function in Eq.28 with the dirty limit coherence length, $\xi\left(T\right)=\sqrt{\pi\hbar D/8k_{B}T}$, and $\alpha$ given by Eq.31, since normalization of the wavefunctions should be independent of the effect of scatterings.
It should also be noted that $\sigma_{xx}^{LV\left(5+7\right)}$ has been derived within the microscopic LV approach while neglecting the $q$ dependence of the renormalized pairing vertex factor (see Eq.5). Inclusion of this dependence would transform Eq. 12 to:
$\sigma_{DOS}^{LV}=2\sigma_{xx}^{LV\left(5+7\right)}$ (32)
$=\left(\frac{e^{2}}{d\hbar}\right)\left(\frac{1}{\pi^{4}}\right)\int d\left(\eta_{\left(2\right)}q^{2}\right)\frac{\psi^{\prime\prime}\left(\frac{1}{2}+\frac{\hbar D}{4\pi k_{B}T}q^{2}\right)}{\varepsilon+\eta_{\left(2\right)}q^{2}}$
In this expression, due to the fact that $\alpha$ is independent of $q$, the CPFs momentum distribution function Eq.28, derived above by neglecting the $q$ dependence of the (Cooperon) factor $\psi^{\prime\prime}\left(1/2+\hbar Dq^{2}/4\pi k_{B}T\right)$, is clearly identified under the integral, separable from any additional $q$ dependent factors, so that $\sigma_{xx}^{LV\left(5+7\right)}$can be written in the momentum dependent Drude-like form:
$\sigma_{xx}^{LV\left(5+7\right)}=-2\frac{e^{2}}{m^{\ast}}\frac{1}{d}\int\frac{d^{2}q}{\left(2\pi\right)^{2}}\left\langle\left|\phi\left(q\right)\right|^{2}\right\rangle\tau_{SO}\left(q\right)$ (33)
Here the additional $q$ dependence appears as an effective correction to the single-electron relaxation time:
$\tau_{SO}\left(q\right)\equiv\tau_{SO}\frac{\psi^{\prime\prime}\left(\frac{1}{2}+\frac{\hbar D}{4\pi k_{B}T}q^{2}\right)}{\psi^{\prime\prime}\left(\frac{1}{2}\right)},$ (34)
$\tau_{SO}\left(q\rightarrow 0\right)=\tau_{SO}$
The resulting expression for $\tau_{SO}\left(q\right)$ shows that the relaxation time of electrons involved in scattering by CPFs (see Fig.2) is effectively suppressed at any exchange of momentum ( $q>0$ ) with a CPF.
Note that Eq.33, without the $q$ dependence of $\tau_{SO}\left(q\right)$ is well defined in the zero temperature limit. The result, after integration over $q$ with the cutoff $q_{c}=1/\xi\left(T\right)$ (see Sec.V for discussion of the cutoff), has a nonvanishing value, that is:
$\sigma_{xx}^{LV\left(5+7\right)}\simeq-\left(\frac{7\zeta\left(3\right)}{\pi^{4}}\right)\left(\frac{e^{2}}{d\hbar}\right)\ln\left(1+\frac{1}{\varepsilon}\right)$ (35)
Taking into account the $q$ dependence of $\tau_{SO}\left(q\right)$ (Eq.34) under the integration over $q$ in Eq.33 significantly suppresses $\sigma_{DOS}^{LV}$ at very low temperature, however, its non vanishing zero temperature limit (see Appendix D) is due to quantum fluctuations.
IV Extension of the microscopic theory to finite magnetic field
Considering, either Eq.1 for the current-current correlator of the AL diagram, or Eq.8 for the correlator of the 5-th LV diagram, the new ingredient associated with the finite magnetic field at low temperatures, besides the important modification introduced to the fluctuation propagator (see below), is the renormalized pairing vertex factor, which includes the effect of spin-orbit scatterings and the magnetic field effect through Zeeman spin splitting $\mu_{B}H=\hslash I$ (see Ref.MZPRB2021 ):
$\lambda_{\pm}\left(q,\omega_{n},-\omega_{n}\right)\equiv\frac{S_{\pm}\left(\omega_{n},q\right)}{S_{\pm}^{0}\left(\omega_{n},q\right)}$ (36)
where:
$S_{\pm}^{0}\left(\omega_{n}\mathbf{,}q\right)\approx\frac{\pi}{\left|\omega_{n}\right|+\frac{1}{2\tau_{SO}}+\frac{1}{2}Dq^{2}\pm iIsign\left(\omega_{n}\right)},$
and:
$S_{\pm}\left(\omega_{n}\mathbf{,}q\right)\approx\pi\frac{\left|\omega_{n}\right|+\frac{1}{\tau_{SO}}+\frac{1}{2}Dq^{2}\mp iIsign\left(\omega_{n}\right)}{\left(\left|\omega_{n}\right|+\frac{1}{2\tau_{SO}}+\frac{1}{2}Dq^{2}\right)^{2}-\left(\frac{1}{2\tau_{SO}}\right)^{2}+I^{2}}$
so that:
$\lambda_{\pm}\left(q,\omega_{n},-\omega_{n}\right)\approx\frac{\left[\left|\omega_{n}\right|+\frac{1}{\tau_{SO}}+\frac{1}{2}Dq^{2}\mp iIsgn\left(\omega_{n}\right)\right]\left[\left|\omega_{n}\right|+\frac{1}{2\tau_{SO}}\pm iIsgn\left(\omega_{n}\right)+\frac{1}{2}Dq^{2}\right]}{\left(\left|\omega_{n}\right|+\frac{1}{2}Dq^{2}\right)\left(\left|\omega_{n}\right|+\frac{1}{\tau_{SO}}+\frac{1}{2}Dq^{2}\right)+I^{2}}$ (37)
For the zero-field case, $I=0$, we recover LV result, Eq.5.
It is seen to correspond to coherent scattering on the same impurity by a pair of free electrons, entering to or emerging from states of fluctuating Cooper-pairs Lopatinetal05 ,GalitLarkinPRB01 .
Another important modification due to the presence of the magnetic field is introduced to the fluctuation propagator. An explicit expression, e.g. for the static propagator derived in Ref.MZPRB2021 , takes the form:
$\mathcal{D}^{H}\left(q\right)=\frac{1}{N_{2D}\Phi\left(q;H\right)}$ (38)
where:
$\Phi\left(q;H\right)=\varepsilon_{H}+a_{+}\left[\psi\left(\frac{1}{2}+f_{-}+x\right)-\psi\left(\frac{1}{2}+f_{-}\right)\right]$
$+a_{-}\left[\psi\left(\frac{1}{2}+f_{+}+x\right)-\psi\left(\frac{1}{2}+f_{+}\right)\right]$ (39)
and $x=\hbar Dq^{2}/4\pi k_{B}T$.  In this expression:
$\varepsilon_{H}\equiv\ln\left(\frac{T}{T_{c0}}\right)+a_{+}\psi\left(\frac{1}{2}+f_{-}\right)+a_{-}\psi\left(\frac{1}{2}+f_{+}\right)-\psi\left(\frac{1}{2}\right)$ (40)
is the critical shift parameter, and:
$a_{\pm}$ $=$ $\frac{1}{2}\left\{1\pm\left[1-\left(\frac{2\mu_{B}H}{\varepsilon_{SO}}\right)^{2}\right]^{-1/2}\right\},$
$f_{\pm}$ $=$ $\frac{\varepsilon_{SO}}{4\pi k_{B}T}\left\{1\pm\left[1-\left(\frac{2\mu_{B}H}{\varepsilon_{SO}}\right)^{2}\right]^{1/2}\right\}$
Eq.38, with Eq.39, for the static fluctuation propagator at finite field, is usually approximated (at small $q$ values), by the linear expansion with respect to the kinetic energy, that is:
$\mathcal{D}^{H}\left(q\right)\simeq\frac{1}{N_{2D}}\frac{1}{\varepsilon_{H}+\eta\left(H\right)\frac{\hbar Dq^{2}}{4\pi k_{B}T}}$ (41)
where:
$\eta\left(H\right)=a_{+}\psi^{\prime}\left(\frac{1}{2}+f_{-}\right)+a_{-}\psi^{\prime}\left(\frac{1}{2}+f_{+}\right)$ (42)
is the field-dependent reduced stiffness coefficient of the fluctuations modes. At zero field, where $\eta\left(H=0\right)=\psi^{\prime}\left(\frac{1}{2}\right)=\pi^{2}/2$, Eq.41 reduces to the static limit of Eq.2. At finite field, $\eta\left(H\right)$ tends to zero with vanishing temperature which effectively leads to complete uniform softening of all fluctuation modes. This remarkable field-induced softening takes place under suppressed pair-breaking condition, $2\mu_{B}H<$ $\hbar/\tau_{SO}$, i.e. within a fields range restricted by the spin-orbit scattering rate. The strong spin-orbit scattering relevant to our paper, i.e. typically with $\tau_{SO}\approx 2\times 10^{-13}\boldsymbol{\mathit{s}}$, is of an order of magnitude rate larger than $\mu_{B}H_{c\parallel 0}/\hbar$. Thus, the softening effect could be realized even in the more usual situations of relatively weak spin-orbit scatterings.
IV.1 The Aslamazov-Larkin paraconductivity diagram
Starting with Eq.1 for the current-current correlator of the AL diagram, following the approximation in which the boson frequency arguments of the effective current vertex $B_{x}\left(q,\Omega_{k},\Omega_{\nu}\right)$ are set to zero, the latter for a finite magnetic field $H$, can be written in the simplified form:
$B_{x}^{H}\left(q,\Omega_{k}=0,\Omega_{\nu}=0\right)\simeq-\frac{1}{2}\pi N_{2D}\frac{1}{2}v_{F}^{2}q_{x}\times$
$k_{B}T\sum_{n}\frac{\lambda_{+}^{H}\left(0,\omega_{n},-\omega_{n}\right)\lambda_{-}^{H}\left(0,\omega_{n},-\omega_{n}\right)}{\left[\left|\widetilde{\omega}_{n}\right|-iIsign\left(\omega_{n}\right)\right]^{3}}$ (43)
and in which we use the $q\rightarrow 0$  limit of the renormalized pairing vertex, Eq.37.
Using the above expression for the effective current vertex, we may write the corresponding correlator in the form:
$Q_{xx}^{AL}\left(i\Omega_{\nu}\right)=-4e^{2}k_{B}T\frac{1}{d}\int\frac{d^{2}q}{\left(2\pi\right)^{2}}$
×$\times\left[B_{x}^{H}\left(q,\Omega_{k}=0,\Omega_{\nu}=0\right)\right]^{2}\Psi\left(q\mathbf{,}\Omega_{\nu}\right)$
where:
$\Psi\left(q\mathbf{,}\Omega_{\nu}\right)=\sum\limits_{k}\mathcal{D}\left(q,\Omega_{\nu}+\Omega_{k}\right)\mathcal{D}\left(q,\Omega_{k}\right)$
Carrying out the boson frequency summation and the analytic continuation: $i\Omega_{\nu}\rightarrow\omega$, we obtain after expansion in small $\omega$, to first order:
$\Psi\left(q\mathbf{,}-i\omega\right)-\Psi\left(q\mathbf{,}0\right)\rightarrow\frac{1}{N_{2D}^{2}}\frac{\frac{i\hbar\omega}{4\pi k_{B}T}\psi^{\prime}\left(\frac{1}{2}+\frac{\left(\mu_{B}H\right)^{2}}{2\pi k_{B}T\varepsilon_{SO}}\right)}{\left[\varepsilon_{H}+\eta\left(H\right)\frac{\hbar Dq^{2}}{4\pi k_{B}T}\right]^{3}}$
in which the linear approximation, Eq.41 for the static fluctuation propagator is used.
Complementing the expression for $\lambda_{\pm}\left(q=0,\omega_{n},-\omega_{n}\right)$ in Eq.37 with the dirty limit condition $\left(\tau_{SO}\mu_{B}H/\hbar\right)^{2}\ll 1$, we find that:
$\lambda_{+}^{H}\left(0,\omega_{n},-\omega_{n}\right)\lambda_{-}^{H}\left(0,\omega_{n},-\omega_{n}\right)\approx\frac{\widetilde{\omega}_{n}^{2}}{\left(\left|\omega_{n}\right|+\tau_{SO}I^{2}\right)^{2}}$ (44)
and the corresponding sheet paraconductivity:
$\sigma_{xx}^{AL}\left(\omega\rightarrow 0\right)=\lim_{\omega\rightarrow 0}\frac{i}{\omega}\left[Q_{xx}^{AL}\left(\omega\right)-Q_{xx}^{AL}\left(0\right)\right]$
$=$ $-4e^{2}k_{B}T\frac{1}{d}\int\frac{d^{2}q}{\left(2\pi\right)^{2}}\left[B_{x}^{H}\left(q,\Omega_{k}=0,\Omega_{\nu}=0\right)\right]^{2}$
×$\times\lim_{\omega\rightarrow 0}\frac{i}{\omega}\left[\Psi\left(q\mathbf{,}-i\omega\right)-\Psi\left(q\mathbf{,}0\right)\right]$
, is written in the form:
$\sigma_{xx}^{AL}\left(\omega\rightarrow 0\right)$ $\simeq$ $e^{2}\frac{1}{2\pi d}\int\frac{d^{2}q}{\left(2\pi\right)^{2}}\left(B_{x}^{H}\left(q,\Omega_{k}=0,\Omega_{\nu}=0\right)\right)^{2}$ (45)
$\times$ $\frac{1}{N_{2D}^{2}}\frac{\hbar\eta_{LV}\left(H\right)}{\left(\varepsilon_{H}+\eta\left(H\right)\frac{\hbar Dq^{2}}{4\pi k_{B}T}\right)^{3}}$
where the effective current vertex is given by:
$B_{x}^{H}\left(q,\Omega_{k}=0,\Omega_{\nu}=0\right)=N_{2D}\frac{D}{2\pi k_{B}T}\eta_{LV}\left(H\right)q_{x}$ (46)
with the (dirty limit) LV version of the reduced stiffness function:
$\eta_{LV}\left(H\right)\simeq\psi^{\prime}\left(\frac{1}{2}+\frac{1}{2}\frac{\left(\mu_{B}H\right)^{2}}{\pi k_{B}T\varepsilon_{SO}}\right)$ (47)
Using Eq.46, Eq.45 reduces to:
[…]
where $\widetilde{x}_{c}=\eta\left(H\right)x_{c},x_{c}=\hbar Dq_{c}^{2}/4\pi k_{B}T$, and $q_{c}$ is the cutoff wavenumber.
It is easy to check (see also Appendix E) that under the dirty limit conditions: $4\pi k_{B}T\tau_{SO}/\hbar,2\mu_{B}H\tau_{SO}/\hbar\ll 1$:
$\eta\left(H\right)\approx\eta_{LV}\left(H\right)\approx\psi^{\prime}\left(\frac{1}{2}+\frac{1}{2}\frac{\left(\mu_{B}H\right)^{2}}{\pi k_{B}T\varepsilon_{SO}}\right)$ (48)
so that finally we find for the AL diagram conductivity at finite field::
$\sigma_{AL}^{mic}\left(H\right)\approx\frac{1}{16}\left(\frac{e^{2}}{d\hbar}\right)\frac{\widetilde{\eta}\left(H\right)}{\varepsilon_{H}\left(1+\frac{\varepsilon_{H}}{\eta\left(H\right)x_{c}}\right)}$ (49)
where:
$\widetilde{\eta}\left(H\right)\equiv\frac{\eta\left(H\right)}{\eta\left(0\right)}=\frac{2}{\pi^{2}}\eta\left(H\right)$ (50)
This result is in full agreement with the AL conductivity obtained within the TDGL functional approach (see in Appendix F an erratum of the original derivation presented in Ref.MZJPC2023 ). This result is expected, of course, in light of the equivalence between the TDGL functional approach and the microscopic LV (diagrammatic) theory applied to the AL paraconductivity calculation, as discussed in detailed in Sec.III.
IV.2 The DOS conductivity diagrams
Starting with Eq.8 for the current-current correlator of the 5-th LV diagram, we use, following LV, the approximate expression for the electronic kernel:
$\Sigma_{xx}^{\left(5\right)H}\left(q\rightarrow 0,\Omega_{k}\rightarrow 0,\Omega_{\nu}\right)=k_{B}T\sum\limits_{n}\lambda_{+}^{H}\left(0,\omega_{n},-\omega_{n}\right)\lambda_{-}^{H}\left(0,\omega_{n},-\omega_{n}\right)I_{xx}^{\left(5\right)}\left(0,\omega_{n},0,\Omega_{\nu}\right)$
with the product of the dirty-limit renormalized pairing vertex functions, Eq.44, arriving at the expression:
$\Sigma_{xx}^{\left(5\right)H}\left(q\rightarrow 0,\Omega_{k}\rightarrow 0,\Omega_{\nu}\right)\rightarrow-\pi N_{2D}v_{F}^{2}\frac{k_{B}T}{\hbar^{3}}\times$
$\left[\left(\sum\limits_{n=-\infty}^{-\nu-1}+\sum\limits_{n=0}^{\infty}\right)\frac{1}{\left(2\left|\omega_{n}\right|+2\tau_{SO}I^{2}\right)^{2}}\frac{1}{\widetilde{\omega}_{n}+\widetilde{\omega}_{n+\nu}}sign\left(\omega_{n}\right)+\sum\limits_{n=-\nu}^{-1}\frac{\left(2\widetilde{\omega}_{n}\right)^{2}}{\left(2\left|\omega_{n}\right|+2\tau_{SO}I^{2}\right)^{2}}\frac{1}{\widetilde{\omega}_{n}+\widetilde{\omega}_{n+\nu}}\left(\frac{1}{\left(\omega_{\nu}+1/\tau\right)^{2}}-\frac{1}{\left(2\widetilde{\omega}_{n}\right)^{2}}\right)\right]$ (51)
This expression is identical to the corresponding result for $\Sigma_{xx}^{\left(5\right)}$ in Ref.LV05 if one replaces in Eq.IV.2 the term $2\tau_{SO}I^{2}$ with $2\times\left(\frac{1}{2}Dq^{2}\right)$. The former term corresponds to the Zeeman energy transfer, $2I=2\mu_{B}H/\hbar$, in the two-electron scattering process occurring at each pairing vertex, whereas the latter corresponds to the kinetic energy transfer close to the Fermi surface, $\mathbf{v}_{F}\cdot\mathbf{q}$ , taking place in such a scattering process.
The combined effect of the Zeeman and kinetic energy transfers on the renormalized pairing vertex factor can be inferred from Eq.37 by rewriting it in an approximate form:
$\lambda_{\pm}^{H}\left(q,\omega_{n},-\omega_{n}\right)\approx\frac{1}{\tau_{SO}}\frac{1}{2\left|\omega_{n}\right|+Dq^{2}+2\tau_{SO}I^{2}}$ (52)
which has been derived under the dirty limit conditions: $1/\tau_{SO}\gg 2\tau_{SO}I^{2},Dq^{2},k_{B}T/\hbar.$ This expression reflects the dual effect of the imbalance kinetic and magnetic energies in removing the zero temperature singularity of the renormalized pairing vertex factor $\lambda_{\pm}^{H}\left(q,\omega_{n},-\omega_{n}\right)$, which takes place equivalently in the orbital and spin spaces.
Performing the fragmented Matsubara frequency summations and the analytic continuation: $\Omega_{\nu}\rightarrow-i\omega$, we find:
$\Sigma_{xx}^{\left(5\right)H}\left(0,0,\Omega_{\nu}\rightarrow-i\omega\right)-\Sigma_{xx}^{\left(5\right)H}\left(0,0,0\right)$
$\approx$ $-i\omega\frac{N_{2D}v_{F}^{2}\tau_{SO}^{2}}{8\pi k_{B}T\hbar}\left[\psi^{\prime}\left(\frac{1}{2}+\frac{\hbar}{4\pi k_{B}T\tau_{SO}}\right)-\frac{3\hbar}{4\pi k_{B}T\tau_{SO}}\psi^{\prime\prime}\left(\frac{1}{2}+\frac{\left(\mu_{B}H\right)^{2}}{2\pi k_{B}T\varepsilon_{SO}}\right)\right]$
in which the first term within the square brackets can be neglected in the dirty limit: $\hbar/4\pi k_{B}T\tau_{SO}\gg 1$.
Adding the complementary contribution of the 7-th diagram, and exploiting the dirty-limit approximation mentioned above, we have for the combined retarded current-current correlator:
$Q_{xx}^{\left(5+7\right)H}\left(i\Omega_{\nu}\rightarrow\omega\right)\rightarrow-i\omega\frac{e^{2}}{d}\frac{D}{4\pi^{2}k_{B}T}\psi^{\prime\prime}\left(\frac{1}{2}+\frac{\left(\mu_{B}H\right)^{2}}{2\pi k_{B}T\varepsilon_{SO}}\right)\int\frac{d^{2}q}{\left(2\pi\right)^{2}}\frac{1}{\varepsilon_{H}+\frac{\eta\left(H\right)\hbar D}{4\pi k_{B}T}q^{2}}$ (54)
Adding the contributions of the 6-th plus 8-th diagrams, which are identical to those of the 5-th and the 7-th ones respectively, we find:
$\sigma_{xx}^{\left(5+6+7+8\right)H}=2\sigma_{xx}^{\left(5+7\right)H}\simeq e^{2}\frac{1}{d\hbar}\frac{1}{\pi^{4}}\psi^{\prime\prime}\left(\frac{1}{2}+\frac{\left(\mu_{B}H\right)^{2}}{2\pi k_{B}T\varepsilon_{SO}}\right)\int d\left(\eta_{\left(2\right)}q^{2}\right)\frac{1}{\varepsilon_{H}+\widetilde{\eta}\left(H\right)\left(\eta_{\left(2\right)}q^{2}\right)}$ (55)
Performing the integration the DOS conductivity at finite field is written as:
$\sigma_{DOS}^{mic}\left(H\right)\simeq-\left(\frac{e^{2}}{2\pi^{2}}\right)\frac{1}{d\hbar}\frac{\left|\psi^{\prime\prime}\left(\frac{1}{2}+\frac{\left(\mu_{B}H\right)^{2}}{2\pi k_{B}T\varepsilon_{SO}}\right)\right|}{\eta\left(H\right)}\ln\left(1+\frac{\eta\left(H\right)x_{c}}{\varepsilon_{H}}\right)$ (56)
with the cutoff: $x_{c}=\hbar Dq_{c}^{2}/4\pi k_{B}T$ ,
V The Cooper-pair fluctuations density at finite field
The detailed analysis presented in Sec.III indicates that Eq.28 for the CPFs momentum distribution function at zero field, derived within the (exclusive boson) TDGL functional approach, can be identified in the microscopic diagrammatic LV theory, while writing the basic $(5+7)$-diagrams contribution to the DOS conductivity in a Drude-like form, Eq.33, with a $q$ dependent single-electron relaxation time $\tau_{SO}\left(q\right)$ (see Eq.34). In the presence of a finite magnetic field, while neglecting the $q$ dependence of the renormalized pairing vertex factor (Eq.52), the microscopic expression for the DOS conductivity, Eq.55, includes a field-dependent Cooperon factor, $\psi^{\prime\prime}\left(1/2+\left(\mu_{B}H\right)^{2}/2\pi k_{B}T\varepsilon_{SO}\right)$, which reflects loss of single electron coherence by the magnetic field, analogous to the corresponding zero-field $q$ dependent factor appearing in Eq.32. In the presence of both the $q$-dispersion and the magnetic field, the loss of coherence originates at the pairing vertices shown in Fig.2 for the DOS conductivity diagram, where pairs of single-electron lines with nonzero center-of-mass momentum ($q\neq 0$) and total magnetic moment ($I\neq 0$) create and then annihilate CPFs. Physically speaking, it is interpreted as suppression of electron coherence-time due to electron scattering with background electrons via virtual exchange of CPFs.
V.1 A proper definition of the Cooper-pair fluctuations density
Repeating the calculation of Sec.IV.B by including the $q$ dependence of the renormalized pairing vertex factor, Eq.52, the result for the DOS conductivity in the presence of magnetic field:
$\sigma_{DOS}^{mic}\left(H\right)=2\sigma_{xx}^{LV\left(5+7\right)}\left(H\right)$ (57)
can be written in a Drude-like form:
$\sigma_{xx}^{LV\left(5+7\right)}\left(H\right)=-2\frac{e^{2}}{m^{\ast}}\frac{1}{d}\int\frac{d^{2}q}{\left(2\pi\right)^{2}}\left\langle\left|\phi\left(\mathbf{q}\right)\right|^{2}\right\rangle_{H}\tau_{SO}\left(q;H\right)$ (58)
which includes both the field and $q$ dependencies of the Cooperon factor, i.e.:
$\tau_{SO}\left(q;H\right)\equiv\frac{\psi^{\prime\prime}\left(\frac{1}{2}+\frac{\hbar D}{4\pi k_{B}T}q^{2}+\frac{\left(\mu_{B}H\right)^{2}}{2\pi k_{B}T\varepsilon_{SO}}\right)}{\psi^{\prime\prime}\left(\frac{1}{2}\right)}\tau_{SO}$ (59)
and the momentum distribution function:
$\left\langle\left|\phi\left(\mathbf{q}\right)\right|^{2}\right\rangle_{H}\equiv\frac{1}{\alpha}\frac{1}{\varepsilon_{H}+\widetilde{\eta}\left(H\right)\left(\eta_{\left(2\right)}q^{2}\right)}$ (60)
with $\alpha$ given by Eq.31.
Similar to the zero-field case discussed in Sec.III (see Ref.LV05 ), the TDGL-Langevin approach will be used below to derive Eq.60 for the momentum distribution function in a magnetic field. The required modification with respect to the zero-field case is associated with the effect of the magnetic field on the fluctuating pairs life-time (given by Eq.27 in the zero-field case). Thus, starting with The TDGL-Langevin equation:
$\widehat{L}^{-1}\phi\left(\mathbf{r},t\right)=\zeta\left(\mathbf{r},t\right)$
under the white-noise condition of the Langevin force correlator:
$\left\langle\zeta^{\ast}\left(\mathbf{r},t\right)\zeta\left(\mathbf{r}^{\prime},0\right)\right\rangle=2k_{B}T\hbar\gamma_{GL}\left(H\right)\delta\left(\mathbf{r}-\mathbf{r}^{\prime}\right)\delta\left(t\right)$ (61)
the (field-dependent) damping parameter $\gamma_{GL}\left(H\right)\equiv\widetilde{\gamma}_{GL}\left(H\right)\pi\alpha/8$ should be determined self consistently with the life time of the fluctuation modes. To determine the TDGL propagator $\widehat{L}$ it would be convenient to go to the wavenumber-frequency representation and exploit its relation (Eq.24) to the microscopically derived dynamical fluctuation propagator (see its static version, Eq.41):
$\mathcal{D}^{H}\left(q,\Omega_{k}\rightarrow-i\Omega\right)\simeq\frac{1}{N_{2D}}\frac{1}{\varepsilon_{H}+\eta\left(H\right)\hbar\frac{Dq^{2}-i\Omega}{4\pi k_{B}T}}$
$=\frac{\alpha k_{B}T}{N_{2D}}L\left(q,\Omega\right)$ (62)
Now, exploiting Eq.61 in evaluating the correlation function, that is:
$\left\langle\phi^{\ast}\left(\mathbf{q};t\right)\phi\left(\mathbf{q};0\right)\right\rangle_{H}=2k_{B}T\hbar\gamma_{GL}\left(H\right)\int\frac{d\Omega}{2\pi}e^{-i\Omega t}\left|L\left(q,\Omega\right)\right|^{2}$
and using Eq. 24 together with Eq.62, the frequency integration is easily performed by the residue method, to find:
$\left\langle\phi^{\ast}\left(\mathbf{q};t\right)\phi\left(\mathbf{q};0\right)\right\rangle_{H}=\frac{\widetilde{\gamma}_{GL}\left(H\right)}{\widetilde{\eta}\left(H\right)}\left(\frac{1}{\alpha}\frac{1}{\varepsilon_{H}+\eta\left(H\right)\frac{\hbar Dq^{2}}{4\pi k_{B}T}}\right)$
×$\times\exp\left[-\frac{\varepsilon\left(q;H\right)}{\widetilde{\eta}\left(H\right)\gamma_{GL}}t/\hbar\right]$ (63)
where the energy of the $q$ fluctuation-mode is given by:
$\varepsilon\left(q;H\right)=k_{B}T\alpha\left(\widetilde{\varepsilon}_{H}+\frac{\eta\left(H\right)\hbar}{4\pi k_{B}T}Dq^{2}\right)$ (64)
At this point one note that in order that the (field-dependent) damping parameter, $\gamma_{GL}\left(H\right)=\widetilde{\gamma}_{GL}\left(H\right)\gamma_{GL}$, of the Langevin force correlator (Eq.61), be determined consistently with the characteristic rate of damping of the correlation function in Eq.63 $\widetilde{\gamma}_{GL}\left(H\right)$ should satisfy the identity:
$\widetilde{\gamma}_{GL}\left(H\right)=\widetilde{\eta}\left(H\right)$ (65)
which reduces Eq.63 to the equivalent of the fluctuation-dissipation theorem, that is:
$\left\langle\phi^{\ast}\left(\mathbf{q};t\right)\phi\left(\mathbf{q};0\right)\right\rangle_{H}=\left\langle\left|\phi\left(\mathbf{q}\right)\right|^{2}\right\rangle_{H}\exp\left(-\frac{t}{\tau_{GL}\left(q;H\right)}\right)$ (66)
where
$\tau_{GL}\left(q;H\right)\equiv\hbar\frac{\widetilde{\eta}\left(H\right)\gamma_{GL}}{\varepsilon\left(q;H\right)}=\hbar\frac{\gamma_{GL}\left(H\right)}{\varepsilon\left(q;H\right)}$ (67)
is the life-time of the fluctuation mode at wavelength $q$. Evidently, as seen in Eq.66, the auto-correlation function, $\left\langle\phi^{\ast}\left(\mathbf{q};0\right)\phi\left(\mathbf{q};0\right)\right\rangle_{H}$ is found to be equal to the equilibrium momentum distribution function, $\left\langle\left|\phi\left(\mathbf{q}\right)\right|^{2}\right\rangle_{H}$ , given by Eq.60.
Returning to the microscopic evaluation of the DOS conductivity, the field dependence, as well as the $q$ dependence, of the Cooperon factor, has different physical origin than that of the remaining integrand factor $\left\langle\left|\phi\left(\mathbf{q}\right)\right|^{2}\right\rangle_{H}$  in Eq.58, which appears as a proper finite-field extension of the zero-field CPFs momentum distribution, Eq.28. Under these circumstances the definition of a field-dependent relaxation time for impurity scattering of electrons, $\tau_{SO}\left(q;H\right)$, in Eq.59 is a natural finite-field extension of Eq.34 for $\tau_{SO}\left(q\right)$ at zero field.
This clear separability between the (collective) fluctuation effect and the single-electron effect in the microscopic theory enables us to define an overall CPFs density:
$n_{S}\left(H\right)\equiv\frac{1}{d}\int\frac{d^{2}q}{\left(2\pi\right)^{2}}\left\langle\left|\phi\left(\mathbf{q}\right)\right|^{2}\right\rangle_{H}$ (68)
$=$ $\left(\frac{7\zeta\left(3\right)E_{F}}{4\pi^{2}k_{B}T}\right)\frac{\pi}{d}\int_{0}^{q_{c}^{2}}\frac{d\left(q^{2}\right)}{\left(2\pi\right)^{2}}\frac{1}{\varepsilon_{H}+\xi^{2}\left(H\right)q^{2}}$
even in the general case where $\tau_{SO}\left(q;H\right)$ depends on $q$ (see Eq.59). Here the field-dependent CPF coherence length is given by:
$\xi\left(H\right)=\sqrt{\widetilde{\eta}\left(H\right)\frac{\pi\hbar D}{8k_{B}T}}$ (69)
The high-field case of interest to us allows, however, to neglect the relatively weak $q$ dependence in Eq.59, so that the DOS conductivity at finite field:
$\sigma_{DOS}^{mic}\left(H\right)=2\sigma_{xx}^{LV\left(5+7\right)}\left(H\right)$ (70)
takes the simple Drude-like form:
$\sigma_{xx}^{LV\left(5+7\right)}\left(H\right)\simeq-2\frac{e^{2}}{m^{\ast}}n_{S}\left(H\right)\tau_{SO}\left(H\right)$ (71)
with:
$\tau_{SO}\left(H\right)\equiv\widetilde{\tau}_{SO}\left(H\right)\tau_{SO}\equiv\frac{\psi^{\prime\prime}\left(\frac{1}{2}+\frac{\left(\mu_{B}H\right)^{2}}{2\pi k_{B}T\varepsilon_{SO}}\right)}{\psi^{\prime\prime}\left(\frac{1}{2}\right)}\tau_{SO}$ (72)
Performing the integration in Eq.68 up to the cut off at $q_{c}^{2}=\xi^{-2}\left(H\right)$ (see Appendix E for further discussion of the cutoff), and substituting in Eq.71, it is written in terms of the basic dimensionless field-dependent parameters $\widetilde{\eta}\left(H\right)$ and $\widetilde{\tau}_{SO}\left(H\right)$, so that:
$\sigma_{DOS}^{mic}\left(H\right)=-\frac{e^{2}}{d\hbar}\left(\frac{14\zeta\left(3\right)}{\pi^{4}}\right)\frac{\widetilde{\tau}_{SO}\left(H\right)}{\widetilde{\eta}\left(H\right)}\ln\left(1+\frac{1}{\varepsilon_{H}}\right)$ (73)
The key parameter in Eq.68, through Eq.69, is the normalized reduced stiffness function $\widetilde{\eta}\left(H\right)$, defined in Eq.50 (see Eq.42), which under the dirty limit conditions, according to Eq.48, can be rewritten in the form:
$\eta\left(H\right)\approx\psi^{\prime}\left(\frac{1}{2}+\frac{T_{H}}{T}\right)$ (74)
where the characteristic temperature $T_{H}$ is defined as:
$T_{H}\equiv\frac{\left(\mu_{B}H\right)^{2}}{2\pi k_{B}\varepsilon_{SO}}$ (75)
(a) Refer to caption (b)
Refer to caption (c)
[…]
At low temperatures and sufficiently high fields, where $T_{H}/T\gg 1$, the asymptotic form of the digamma function yields vanishing stiffness at any finite field, i.e.:
$\eta\left(H\right)\rightarrow\frac{T}{T_{H}},T\ll T_{H}$ (76)
which remains finite, however, (equal to $\eta\left(0\right)=\pi^{2}/2$) at zero field.
This extreme low temperature softening of the fluctuation modes at finite field results in divergent CPFs density (see Fig.3 and more details in Appendix E):
$n_{S}\left(H\right)\rightarrow n_{S}^{0}\left(\frac{T_{H}}{T}\right)\ln\left(1+\frac{1}{\varepsilon_{H}}\right),T\ll T_{H}$ (77)
where the field and temperature independent CPFs density parameter $n_{S}^{0}$ is defined as:
$n_{S}^{0}\equiv\left(\frac{7\zeta\left(3\right)E_{F}}{4\pi^{2}\hbar D}\right)\frac{1}{d}=\frac{7\zeta\left(3\right)}{2\pi}\frac{1}{d}N_{2D}\varepsilon_{SO}$ (78)
This divergence is, of course unphysical, since $n_{S}\left(H\right)$ could not exceed twice the normal-state electron density $n_{0}=k_{F}^{2}/2\pi$, reflecting the neglect of quantum tunneling of CPFs in our model at zero temperature (see more details in Refs. MZPRB2021 ,MZJPC2023 and in the discussion section below).
The corresponding limiting coherence length is finite, diminishing with increasing magnetic field:
$\xi\left(H\right)\rightarrow\frac{1}{\mu_{B}H}\sqrt{\varepsilon_{SO}\hbar D/2}=\frac{\hbar v_{F}}{2\mu_{B}H},T\ll T_{H}$ (79)
In contrast, at zero field, the reduced stiffness function remains finite at any temperature (i.e. $\widetilde{\eta}\left(H=0\right)=1$), and the resulting CPFs density is also finite, equal to:
$n_{S}\left(H=0\right)=n_{S}^{0}\left(\frac{2}{\pi^{2}}\right)\ln\left(1+\frac{1}{\varepsilon}\right)$ (80)
but with infinitely long coherence length for $T\rightarrow 0$:
$\xi\left(H=0\right)\rightarrow\frac{1}{4}\hbar v_{F}\sqrt{\frac{\pi}{k_{B}T\varepsilon_{SO}}},T\ll T_{H}$ (81)
It will be instructive at this point to use in Eq.68 an extension of Eq.60 for the momentum distribution function to large wavenumbers, which can be derived in the dirty limit, $\hbar/\tau_{SO}=\varepsilon_{SO}>>\mu_{B}H,k_{B}T$, i.e. (see Appendix E):
$\left\langle\left|\phi\left(\mathbf{q}\right)\right|^{2}\right\rangle_{H}=\frac{7\zeta\left(3\right)E_{F}}{4\pi^{2}k_{B}T}\frac{1}{\Phi\left(q;H\right)}$ (82)
with:
$\Phi\left(q;H\right)=\widetilde{\varepsilon}_{H}+\psi\left(\frac{1}{2}+\frac{T_{c0}}{T}\frac{2}{\pi^{2}}\xi_{0}^{2}q^{2}+\frac{T_{H}}{T}\right)-\psi\left(\frac{1}{2}+\frac{T_{H}}{T}\right)$ (83)
and (see Eq.26):
$\xi_{0}\equiv\xi\left(T_{c0}\right)=\sqrt{\frac{\hbar\pi D}{8k_{B}T_{c0}}}$ (84)
Note the decorated notation $\widetilde{\varepsilon}_{H}$ used explicitly in Eq.83 for the positive-definite critical shift parameter, which was introduced originally in Ref. MZPRB2021 to take into account, self-consistently, the effect of interaction between fluctuations on $\varepsilon_{H}$ defined in Eq.40 for free (Gaussian) fluctuations. We shall use this notation for the critical shift parameter from now on in this paper to remind the reader that it can have only nonvanishing positive values, except for $T=0$, consistently with the absence of zero resistance in the experimental data.
The importance of the vanishing reduced stiffness function, $\widetilde{\eta}\left(H\right)\simeq\left(2/\pi^{2}\right)\psi^{\prime}\left(1/2+T_{H}/T\right)\rightarrow\left(2/\pi^{2}\right)\left(T/T_{H}\right)$, of the fluctuation modes at low temperatures, $\left(T/T_{H}\right)\ll 1$, becomes transparent only after expanding the energy function, Eq.83, to first order in the kinetic energy term, which yields in agreement with Eq.60: $\Phi\left(q;H\right)\simeq\widetilde{\varepsilon}_{H}+\widetilde{\eta}\left(H\right)\left(T_{c0}/T\right)\xi_{0}^{2}q^{2}=\widetilde{\varepsilon}_{H}+\widetilde{\eta}\left(H\right)\left(\eta_{\left(2\right)}q^{2}\right)$ (see Appendix E). Continuing analytically the Taylor expansion of $\Phi\left(q;H\right)$ in $\widetilde{\eta}\left(H\right)\left(T_{c0}/T\right)\xi_{0}^{2}q^{2}\rightarrow\left(2/\pi^{2}\right)\left(T_{c0}/T_{H}\right)\xi_{0}^{2}q^{2}$, it takes the low temperature form (see Appendix E):
$\Phi\left(q;H\right)\rightarrow\widetilde{\varepsilon}_{H}+\ln\left[1+\left(\frac{2}{\pi^{2}}\right)\left(\frac{T_{c0}}{T_{H}}\right)\xi_{0}^{2}q^{2}\right]$ (85)
in which the logarithmic term does not depend on temperature ! The condition for the validity of the linear approximation, uniformly for all $q$ values up to the cutoff $q_{c}$:
$\left(\frac{T_{c0}}{T_{H}}\right)x_{c0}\ll 1$ (86)
where:
$x_{c0}\equiv\frac{\hbar Dq_{c}^{2}}{4\pi k_{B}T_{c0}}=\frac{2}{\pi^{2}}\xi_{0}^{2}q_{c}^{2}$ (87)
depends both on the field, through $T_{H}$, and on the cutoff wavenumber, through $x_{c0}$ (see Fig.3), but does not depend on temperature.
It is remarkable that even for larger values of the cutoff parameter $x_{c0}$, where the linear approximation (see Eq.86) is not valid, the low temperature (asymptotic) CPFs density, given by:
$n_{S}\left(H\right)\rightarrow n_{S}^{0}\left(\frac{T_{H}}{T}\right)\int_{0}^{x_{c0}T_{c0}/T_{H}}\frac{d\chi}{\widetilde{\varepsilon}_{H}+\ln\left(1+\chi\right)}$ (88)
has the same $1/T$ divergence at $T\ll T_{H}$ as the linear approximation given by Eq.77, but with a larger slope, as shown in Fig.3 (see also Appendix E). It should be emphasized that this divergence is not peculiar to the strong spin-orbit scattering situation considered in the present paper. As indicated in Sec.IV, the field-induced ultimate softening of the fluctuation modes responsible for this divergence takes place at sufficiently low temperature at any field that induces Zeeman spin-splitting (pair-breaking) frequency, $2\mu_{B}H/\hbar$, smaller than the (pair-compensating) spin-orbit relaxation rate $1/\tau_{SO}$.
V.2 The crossover to localization and phase separation of Cooper-pair fluctuations
The discontinuous nature of the reduced stiffness function $\eta\left(H\right)$ at $H=0$ in the zero-temperature limit, revealed in Eq. 76, is inherent to peculiar features of the CPFs at low temperatures. These features may be best appreciated by considering the Fourier transform to real space of the momentum distribution function, Eq.60, in the low temperature limit $T\ll T_{H}$, where field-induced condensation in real space is expected on the basis of the limiting behavior of $n_{S}\left(H\right)$ and $\xi\left(H\right)$ according to Eqs.77 and 79, respectively. In the absence of interaction between fluctuations this Fourier transform can be shown to be related, by Eq.66, to the equal-time real-space Cooper-pair amplitude correlation function through the equation:
$\left\langle\phi^{\ast}\left(\mathbf{r+\rho,}t_{0}\right)\phi\left(\mathbf{r,}t_{0}\right)\right\rangle_{H}=\left(\frac{1}{2\pi}\right)^{2}\int d^{2}q\left\langle\left|\phi\left(\mathbf{q}\right)\right|^{2}\right\rangle_{H}e^{i\mathbf{q\cdot\rho}}$
$\equiv\left\langle\left|\phi\left(\mathbf{\rho}\right)\right|^{2}\right\rangle_{H}$ (89)
The last identity indicates that, as a function of $\mathbf{\rho}$, this correlation function is a real space measure of the CPFs density, which determines the probability amplitude for a CPF, generated at any point $\mathbf{r}$ of the underlying uniform 2D system, to propagate a distance $\rho=\left|\mathbf{\rho}\right|$.
Performing the angular integration in Eq.89 and using the more general expression for the momentum distribution function in the dirty limit, Eq.82, i.e. with $\Phi\left(q;H\right)$ given by Eq.83, it can be rewritten in the form:
$\left\langle\left|\phi\left(\mathbf{\rho}\right)\right|^{2}\right\rangle_{H}=\frac{1}{2\pi}n_{S}^{0}\xi_{0}^{-2}g\left(\rho/\xi_{0}\right)$ (90)
where $g\left(\rho/\xi_{0}\right)$ is given by:
[…]
$\widetilde{q}\equiv\xi_{0}q$, and $J_{0}\left(\widetilde{q}\rho/\xi_{0}\right)$ is the zero order Bessel function of the first kind.
The dependence of the dimensionless density function $g\left(\rho/\xi_{0}\right)$ on $\rho$ has a decaying envelope, modulated by an oscillatory function associated with the sharp cutoff $q_{c}$. The length scale of this attenuation is the localization length, given by (see Appendix G):
$\rho_{loc}\left(H\right)=\left(\frac{T_{c0}}{T}\frac{\widetilde{\eta}\left(H\right)}{\widetilde{\varepsilon}_{H}}\right)^{1/2}\xi_{0}$ (92)
The results of a detailed analysis of $g\left(\rho/\xi_{0}\right)$ and its asymptotic behavior (see Appendix G) are shown in Fig.4, where the condensing CPFs at diminishing temperature are found to pile up within a region of diminishing size of the order of $\rho_{loc}\left(H\right)$ under increasing field $H$.
The positive-definite self-consistent critical shift parameter, $\widetilde{\varepsilon}_{H}$, appearing in Eq.92, which is a monotonically increasing function of the field $H$, is used in the calculations as an independent free parameter, in order to illustrate how the minimal gap, $\widetilde{\varepsilon}_{H}$, of the energy spectrum given in Eq.83, influences the condensation and localization of CPFs in real space.
(a) Refer to caption (b)
Refer to caption (c) Refer to caption (d)
[…]
The CPFs density function defined in Eq.90, which shows localization of fluctuations propagated around any arbitrary point in the 2D system, has been derived with the assumption that interactions between fluctuations are neglected. Thus, it is expected that Coulomb repulsion between the fluctuations would lead to formation of an extended 2D structure of mesoscopic puddles. This is qualitatively plausible conclusion, though we have not gone beyond that point quantitatively since the details of this structure is not relevant to our purposes here (see the remark at the end of Sec.VI).
One should also note that the validity of Eq.90 as a local density of some complex bosonic particles as a function of the distance from their point of creation is limited by the life-time $\tau_{GL}\left(q;H\right)$ of the corresponding excitation, as given by Eq. 67. At low temperatures, $T\ll T_{H}$, where:
$\tau_{GL}\left(q;H\right)\rightarrow\frac{\hbar/4\pi k_{B}T_{H}}{\widetilde{\varepsilon}_{H}+\hbar Dq^{2}/4\pi k_{B}T_{H}}$
our estimate of this characteristic time at $H=4T$, and for the typical experimental parameters (for which $T_{H}=33mK$, and the cutoff parameter $x_{c0}\approx 0.015$), is: $\tau_{GL}\left(q_{c};H=4T\right)\approx 10^{-10}\boldsymbol{\mathit{s}}$, that is about three orders of magnitude larger than the typical electron relaxation time $\tau_{SO}=\hbar/\varepsilon_{SO}\approx 10^{-13}\boldsymbol{\mathit{s}}$. Within this relatively long time scale our theory predicts enormously large accumulated charging of small (mesoscopic) spatial regions (puddles) by CPFs upon increasing magnetic field, which due to Coulomb repulsion, should push out unpaired normal-state electrons, leading eventually to phase separation between the two charged subsystems. Under these circumstances the normal-state electrons are not directly affected by CPFs, but are only holding with them mutual charge exchange under interphase dynamical equilibrium. In particular, regarding the transport problem under study here, the resulting normal-state electron gas keeps its relaxation time $\tau_{SO}$ unaffected by CPFs (see Sec.VI).
To summarize, upon increasing magnetic field, the system of CPFs crossovers from a spatially uniform state, at zero field, to inhomogeneous states of condensed mesoscopic puddels, at high fields. For sufficiently low temperatures and high fields, where $T\ll T_{H}$, the inhomogeneous states are characterized by enormously large local charging due to CPFs condensation in mesoscopic puddles, separated from the unpaired normal-state electrons in inter-puddel regimes.
A plausible estimate of the crossover field, $H_{cross}$, can be made on the scale of the reduced stiffness parameter $\widetilde{\eta}\left(H\right)$, half way between its edge values corresponding the uniform and the nonuniform states, that is:
$\widetilde{\eta}\left(H_{cross}\right)=1/2$
At low temperatures, where $T\ll T_{H}$, this may be done by using Eq.76, which yields:
$\mu_{B}H_{cross}\sim\left(\varepsilon_{SO}k_{B}T\right)^{1/2}$ (93)
For the sake of better sense of scaling, we may estimate the field, $H_{upp}$ at which the low-temperature limit of the CPFs density $n_{S}\left(H\right)$, Eq.77, approaches its (physical) upper limit, i.e. $1/2$ of the total number density of electrons $n_{0}=k_{F}^{2}/2\pi d$:
$n_{S}\left(H\right)$$\rightarrow$ $\left(\frac{7\zeta\left(3\right)}{4\pi^{2}}\right)\frac{1}{d}N_{2D}\frac{\left(\mu_{B}H\right)^{2}}{k_{B}T}\ln\left(1+\frac{1}{\widetilde{\varepsilon}_{H}}\right)$
$=$ $\frac{1}{2}\left(\frac{1}{d}\frac{k_{F}^{2}}{2\pi}\right)=\frac{1}{2}n_{0}$
which yields:
$\mu_{B}H_{upp}\sim\left(E_{F}k_{B}T\right)^{1/2}$ (94)
Thus, the plausible hierarchy of the parameters $E_{F}>\varepsilon_{SO}$ ensures that the crossover field, $H_{cross}$ can always be reached.
VI Discussion and Conclusion
The peculiar features of the system of CPFs, discussed in the previous sections, reflect essential inconsistency of the microscopic theory of fluctuations in superconductors at finite field and very low temperatures. In particular, it was shown that the contribution of the basic pair of diagrams to the DOS conductivity obtained within the conventional microscopic approach can be expressed in terms of an effective Drude formula (see Eq.71) in which $n_{S}\left(H\right)$ is an effective CPFs density (Eq.68) and $\tau_{SO}\left(H\right)$ is an effective single electron relaxation time (Eq.72). These definitions are not arbitrary; on the one hand, the field dependence of $\tau_{SO}\left(H\right)$ is exclusively determined by the renormalized pairing vertex factors (see Eq.52), associated with electron scatterings by background electrons via virtual exchange of CPFs. On the other hand, the field dependence of $n_{S}\left(H\right)$ is associated exclusively with the fluctuation propagator, controlled by the normalized reduced stiffness function $\widetilde{\eta}\left(H\right)$, which tends to zero at any $H>0$ with $T\rightarrow 0$ (see Eq.76), but remains finite (equal to $1$) at zero field. Consequently, the density $n_{S}\left(H\right)$ diverges for $T\rightarrow 0$ at any $H>0$ (see Eq.77 and note the comment following Eq.78), but remains finite at $H=0$ (see Eq.80). Furthermore, the divergent Cooper-pair coherence length $\xi\left(H=0\right)$ in the $T\rightarrow 0$ limit (Eq.81), which characterizes a homogeneous system with long-ranged superconducting order at $H=0$, contracts at $H>0$ to a finite length of localization, where increasingly large numbers of CPFs condense under diminishing temperature.
The state-of-the-art microscopic theory of fluctuations in superconductors which leads to these peculiar results is basically a perturbation theory, developed for spatially homogenous systems, which rests upon a diagrammatic expansion of the conductivity in the fluctuations propagator about the normal-state conductivity. In this framework the scope of the calculations is restricted to fluctuations effect on the normal-state conductivity above the transition to superconductivity. Within this microscopic approach we may write the total conductivity as:
$\sigma_{tot}^{mic}\left(H\right)=\sigma_{n}+\sigma_{DOS}^{mic}\left(H\right)+\sigma_{AL}^{mic}\left(H\right)$ (95)
where $\sigma_{n}$ is the normal-state (zero-order in the expansion) conductivity, $\sigma_{DOS}^{mic}\left(H\right)$ is the DOS conductivity given by Eq.73, and $\sigma_{AL}^{mic}\left(H\right)$ is the AL conductivity written in Eq.49. As indicated in Sec.III, the Maki-Thompson conductivity (MakiPTP1968 ,ThompsonPRB1970 ) is neglected here due to the presence of strong spin-orbit scatterings (see Appendix A). Both $\sigma_{DOS}^{mic}\left(H\right)$ and $\sigma_{AL}^{mic}\left(H\right)$ have been calculated here within the framework of the LV method used in Ref.LV05 in which quantum critical fluctuations were neglected.
Under these circumstances at finite field $H$ and very low temperature $T\ll T_{H}$, due to the vanishing $\widetilde{\eta}\left(H\right)\rightarrow\left(2/\pi^{2}\right)T/T_{H}$ and the more quickly vanishing $\widetilde{\tau}_{SO}\left(H\right)\rightarrow\left(T/T_{H}\right)^{2}/7\zeta\left(3\right)$, both the AL (Eq.49) and DOS (Eq.73) conductivities:
$\sigma_{AL}^{mic}\left(H\right)$$\rightarrow$ $\frac{e^{2}}{8\pi^{2}d\hbar}\left(\frac{T}{T_{H}}\right)\frac{1}{\widetilde{\varepsilon}_{H}},$ (96)
$\sigma_{DOS}^{mic}\left(H\right)$$\rightarrow$ $-\frac{e^{2}}{2\pi^{2}d\hbar}\left(\frac{T}{T_{H}}\right)\ln\left(\frac{1}{\widetilde{\varepsilon}_{H}}\right)$ (97)
vanish with $T$.  This linearly vanishing with temperature AL and DOS conductivities are consistent with the results reported in Ref.Glatzetal2011 at very low temperatures in the absence of quantum critical fluctuations (see a remark below Eq.10). The different field dependencies should be related to the different magnetic field orientations (perpendicular in Ref.Glatzetal2011 as compared to parallel in our case).
The above discussion clearly indicates that this perturbation theory can not be directly applied to the extremely inhomogeneous real-space Cooper-pairs (boson) condensate that emerges from our analysis. In particular, although the leading-order (i.e. the DOS and the AL) diagrams contribution to the fluctuation conductivity vanishes for $H>0,T\rightarrow 0$, the definition of the (zeroth order) normal-state conductivity should be drastically modified by the (divergent) CPFs density $n_{S}\left(H\right)$ through their charge exchanges with the (unpaired) normal-state electrons. Thus, formally speaking, while using the simple Drude expression $\left(e^{2}/m^{\ast}\right)n_{0}\tau_{n}$ for $\sigma_{n}$ in Eq.95 with $\tau_{n}$– the normal-state electron relaxation time (equal in our model to $\tau_{SO}$), one notes that it is valid only under the usual situation of nearly uniform real-space distribution of CPFs. However, under the unusually inhomogeneous situation, outlined above, of divergent number density $n_{S}\left(H\right)$ of CPFs in mesoscopic puddles, their dynamical equilibrium with the normal-state electrons should lead, under the assumption of a fixed total number of electrons, to the following modification:
$\sigma_{n}\rightarrow\sigma_{n}^{Mod}\left(H\right)=\frac{e^{2}}{m^{\ast}}\left(n_{0}-2n_{S}\left(H\right)\right)\tau_{SO}$ (98)
$=$ $\frac{e^{2}}{m^{\ast}}n_{0}\tau_{SO}-\frac{e^{2}}{m^{\ast}}2n_{S}\left(H\right)\tau_{SO}$
The total consumption of electrons in the process of pairing sets an upper bound on $n_{S}\left(H\right)$, i.e. $2n_{S}\left(H\right)\leq n_{0}$, corresponding to the ultimate vanishing of $\sigma_{n}\left(H\right)$ for $H>0,T\rightarrow 0$, that is physically equivalent to infinite high-field MR. This ultimate vanishing of both the fluctuation and the normal-state conductivities in the zero temperature limit is due to our neglect of quantum fluctuations. As discussed in detail in several recent papers (Glatzetal2011 , Lopatinetal05 , Khodas2012 ), corrections due to quantum fluctuations in both the DOS and the AL conductivities (as well as in Maki-Thompson contributions, neglected here, see Appendix A) around the quantum critical field prevent the vanishing of the fluctuation conductivity in the zero temperature limit. Furthermore, as discussed in Ref.MZPRB2021 ,MZJPC2023 , quantum tunneling fluctuations (see below) suppress the divergent CPFs density and so prevent the vanishing of $\sigma_{n}^{Mod}\left(H\right)$ in Eq.98.
In the less extreme situations of finite fields and low temperatures, above the crossover field, Eq.93, the second term on the RHS of Eq.98 is identified with the DOS conductivity derived within our TDGL functional approach. Thus, as long as $\widetilde{\tau}_{SO}\left(H\right)\ll 1$, the result of replacing $\sigma_{n}$ in Eq.95 with $\sigma_{n}^{Mod}\left(H\right)$ is seen to agree with the expression for the total conductivity derived within our TDGL functional approach MZJPC2023 .
The crucial point here is in the existence of condensed mesoscopic puddles of long-lived boson excitations, which act through interphase pair-breaking processes as reservoirs for the phase-separated gas of (unpaired) conduction electrons. Within this picture the conductance of the boson excitations is exclusively represented by the Schmidt-Fulde-Maki paraconductivity (see Eq.15), shown above to be equivalent to the Aslamozov-Larkin conductivity in the microscopic diagrammatic approach. The remaining system of unpaired electrons, restricted to regions of diminishing scattering by CPFs, keeps its normal-state relaxation time $\tau_{SO}$, so that its conductance is influenced by the Cooper-pair bosons only through interphase charge exchange, as expressed in Eq.98.
The resulting expression for the total conductance, for $T\ll T_{H}$, $H\geq H_{cross}$:
$\sigma_{tot}\left(H\right)\rightarrow\sigma_{n}\left(H\right)-2\frac{e^{2}}{m^{\ast}}n_{S}\left(H\right)\tau_{SO}+\sigma_{AL}\left(H\right)$ (99)
includes a well defined normal-state conductivity, $\sigma_{n}\left(H\right)$, which is now completely independent of the pairing effect. Its field dependence is introduced to account for possible normal-state MR effects, such as weak localization etc…Hikami-etal-PTP1980 , Fukuyama1985 .
Eq.99 is the main result of physical content in the present paper. Its validity is bound to a complete phase separation between the negatively charged clouds of the CPFs puddles and the unpaired electrons, which can be realized at sufficiently high field ($H\geq H_{cross}$) and low temperature ($T\ll T_{H}$). The picture of charged boson condensates, phase separated from normal-electron regions, which hold mutual dynamical equilibrium via charge exchange, leads one naturally to expect that quantum tunneling of CPFs should take place across GL energy barriers, built-up on puddles’ boundaries. Of course, a certain interphase pair-breaking process sustains the charge exchange across the barriers. The combined quantum-tunneling-pair-breaking (QTPB) process should be, therefore, responsible for suppressing the divergent $n_{S}\left(H\right)$ at $T\rightarrow 0$, under the condition of total charge conservation, such that:
$\lim_{T\rightarrow 0}2n_{S}^{QTPB}\left(H\right)\leq n_{0}\left(H\right)=\left(m^{\ast}/\tau_{SO}e^{2}\right)\sigma_{n}\left(H\right)$
We have presented in our previous papers MZPRB2021 ,MZJPC2023 a phenomenological approach to account for these quantum processes. Their underlying microscopic quantum dynamics remains a demanding challenge in the GL theory.
The influence of quantum critical fluctuations on the AL and DOS conductivities at finite field in the zero temperature limit has been evaluated within the conventional microscopic theory, both for parallel (Lopatinetal05 , Khodas2012 ) and perpendicular (Glatzetal2011 ) field orientations. Typically, it has been found to introduce corrections to the vanishing ”classical” results of magnitude similar to (see e.g. Glatzetal2011 ):
$\delta\sigma_{xx}^{Q}\left(H\right)=-\frac{2e^{2}}{3\pi^{2}d\hbar}\ln\left(\frac{1}{\widetilde{\varepsilon}_{H}}\right)$ (100)
These corrections have been found, however, to be too small for sufficiently influencing the normal-state conductivity in any attempt to account for the pronounced MR peaks observed experimentally in Ref.Mograbi19 . In contrast, at sufficiently low temperatures $T\ll T_{H_{cross}}$, the correction to the normal state conductivity in Eq.99, due to charge transfer to CPFs, which has an effective DOS conductivity form with $\widetilde{\tau}_{SO}\left(H=0\right)=1$, that is: $\sigma_{DOS}^{eff}\left(H\right)\equiv-2\left(e^{2}/m^{\ast}\right)n_{S}\left(H\right)\tau_{SO}$ , yields:
$\sigma_{DOS}^{eff}\left(H\right)\sim-\frac{3.5\zeta\left(3\right)e^{2}}{\pi^{2}d\hbar}\left(\frac{T_{H}}{T}\right)\ln\left(\frac{1}{\widetilde{\varepsilon}_{H}}\right)$ (101)
This result is, in principle, capable of producing the observed large MR peaks, their actual intensities depend, however, on the quantum tunneling effect discussed above.
In conclusion, promoting the hypothesis that the large MR observed at low temperatures, just above a crossover field to superconductivity, in the electron-doped SrTiO3/LaAlO3 interfaces Mograbi19 ,RoutPRL2017 , is associated with CPFs, it is shown that the state-of-the-art microscopic theory of fluctuations in superconductors fails to account for this remarkable effect. This failure is attributed here to the restrictive nature of this theory which is limited to expansions about homogeneous electron systems.
A proper definition of the density of CPFs within this microscopic theory shows, however, that at sufficiently low temperatures the negatively charged CPFs under increasing field tend to condense within mesoscopic puddles so that, the underlying (unpaired) normal-state electron gas should be pushed by Coulomb repulsion away from the puddles territory. The strongly enhanced MR therefore arises from charge transfer, under dynamical equilibrium, between rarefying, normal-state electron gas and condensing CPFs, localized in mesoscopic puddles, separately from the unpaired electrons.
It should be stressed that Coulomb repulsion is not included explicitly in our microscopic theory. However, the plausible assumption of interphase dynamical equilibrium of charges between CPFs and unpaired electrons enables us to arrive at the main result of physical content in this paper, i.e. Eq.99, without a detailed quantitative theory of the underlying complex mechanisms of Cooper-pairing and electron-electron interactions.
Appendix A The relevant diagrams
The literature dealing with the effect of superconducting fluctuations on the conductivity from the point of view of the microscopic GGL theory is quite extensive. The most comprehensive and elaborated account of this approach can be found in Ref.LV05 . Some specific details relevant to our analysis of the diagrams shown in Fig.1, concerning in particular, the relative importance of the various contributions to the DOS conductivity (diagrams 5-8 in Fig.1), can be found in earlier papers: AltshulerJETP1983 , Dorin-etal-PRB1993 .
A different type of diagrams (2-4 in Fig.1), corresponding to electron-hole (Andreev-like) scatterings by CPFs ManivAlexanderJPC1976 , that is well-known as the Maki–Thompson diagram MakiPTP1968 ,ThompsonPRB1970 , includes two parts of contributions: A singular part, arising from the coherent Andreev-like scattering, with positive contribution to the fluctuation conductivity, and a regular part with negative contribution to the conductivity, similar to the DOS conductivity. In our model we disregard this type of diagrams altogether since the strong spin-orbit scatterings, which characterize the SrTiO3/LaAlO3 interfaces under consideration here, are known to destroy the coherence responsible for the singularity LV05 , so that the remaining regular positive contribution is cancelled, or nearly cancelled by the negative ones ManivAlexanderJPC1977 .
Appendix B The Copper-pairs (GL) current density
Eq.13 is obtained from the variational condition of the electromagnetically modified GL free energy functional with respect to the vector potential:
$\frac{\partial\mathfrak{L}\left(\Delta,\mathbf{A}\right)}{\partial\mathbf{A}}+\frac{1}{8\pi}\frac{\partial}{\partial\mathbf{A}}\int d^{3}r\left(\mathbf{\nabla}\times\mathbf{A}\right)^{2}=0$ (102)
in conjunction with the identity:
$\frac{1}{8\pi}\frac{\partial}{\partial\mathbf{A}}\int d^{3}r\left(\mathbf{\nabla}\times\mathbf{A}\right)^{2}=\frac{1}{4\pi}\mathbf{\nabla}\times\left(\mathbf{\nabla}\times\mathbf{A}\right)=\frac{1}{c}\mathbf{j}$
Appendix C The normalization constant
To find $\alpha$ we note that Eq.24, relating the propagators, $\mathcal{D}\left(q,\Omega\right)$ in the $\Delta$ representation to $L\left(q,\Omega\right)$ in the $\phi$ representation, is equivalent to the normalization of the GL wavefunction:
$\left|\phi\left(\mathbf{r}\right)\right|^{2}=\frac{N_{2D}}{\left(\alpha k_{B}T\right)}\left|\Delta\left(\mathbf{r}\right)\right|^{2}$ (103)
As emphasized in the main text, the normalization constant $\alpha$, in the dirty-limit under study, may be evaluated in the clean limit, i.e. for coherence length:
$\xi\left(T\right)\rightarrow\xi_{c}\left(T\right)=\sqrt{\frac{7\zeta\left(3\right)}{8}}\left(\frac{\hbar v_{F}}{2\pi k_{B}T}\right)$ (104)
since normalization of the wavefunctions should not depend on scatterings. Thus, using the expression for $\alpha$ presented in Eq.31, for which the normalization takes the form:
$\left|\phi\left(\mathbf{r}\right)\right|^{2}=\frac{7\zeta\left(3\right)}{8}\frac{k_{F}^{2}}{2\pi}\left(\frac{\left|\Delta\left(\mathbf{r}\right)\right|}{\pi k_{B}T}\right)^{2}\simeq\frac{k_{F}^{2}}{2\pi}\left(\frac{\left|\Delta\left(\mathbf{r}\right)\right|}{\pi k_{B}T}\right)^{2}$ (105)
the clean-limit GL propagator, Eq.25, at zero frequency, is written in the canonical Schrodinger-like form:
$L\left(q,0\right)^{-1}=\frac{\hbar^{2}}{4m^{\ast}}\left[\xi\left(T\right)^{-2}\varepsilon+q^{2}\right]$ (106)
with the Cooper-pair mass equals twice the electron band mass $m^{\ast}$.
The dirty-limit momentum distribution function, Eq.28, may be therefore evaluated with $\alpha$ given by Eq.31 with the dirty-limit coherence length given in Eq.26.  The result takes the form:
$\left\langle\left|\phi\left(q\right)\right|^{2}\right\rangle=\frac{28\zeta\left(3\right)}{\pi^{2}}\left(N_{2D}\frac{\hbar}{\tau_{SO}}\right)\frac{1}{\xi\left(T\right)^{-2}\varepsilon+q^{2}}$ (107)
with the dirty limit coherence length:
$\xi\left(T\right)\rightarrow\xi_{d}\left(T\right)=\sqrt{\frac{\pi\hbar D}{8k_{B}T}}$ (108)
For comparison, the clean limit result is obtained from Eq.107 by replacing $\hbar/\tau_{SO}$ with $\left[4\pi^{3}/7\zeta\left(3\right)\right]k_{B}T$ and $\xi_{d}\left(T\right)$ with $\xi_{c}\left(T\right)$ for the coherence length $\xi\left(T\right)$.
Finally, the TDGL expression for the DOS conductivity, defined in terms of the CPFs density $n_{s}$ by using using Eq.107, is shown here to coincide with the basic microscopic (diagrammatic) result, that is:
$\sigma_{DOS}^{TDGL}\equiv-2n_{s}\frac{e^{2}}{m^{\ast}}\tau_{SO}\equiv$
$-\frac{e^{2}}{\hbar d}\frac{7\zeta\left(3\right)}{\pi^{4}}\int\frac{d\left(\xi\left(T\right)^{2}q^{2}\right)}{\varepsilon+\xi\left(T\right)^{2}q^{2}}=\sigma_{xx}^{LV\left(5+7\right)}$ (109)
Appendix D The dispersion suppressed zero-field DOS conductivity
Consider the corrected zero-field DOS conductivity, Eq.33 , which includes the $q$ dependence of the single-electron relaxation time $\tau_{SO}\left(q\right)$, that is:
[…]
where $x=\pi\hbar Dq^{2}/8k_{B}T$, and: $\psi^{\prime\prime}\left(1/2+2x/\pi^{2}\right)=-2\sum\limits_{n=0}^{\infty}\left(n+1/2+2x/\pi^{2}\right)^{-3}$.  The result is nonvanishing negative function of $\varepsilon$ ($0<\varepsilon\ll 1$), which may be estimated, using two-parameter fitting scheme, to be:
$\sigma_{DOS}^{LV,corr}\approx-\left(\frac{1}{2\pi^{4}}\right)\left(\frac{e^{2}}{d\hbar}\right)\left[7\ln\left(1+\frac{1}{\varepsilon}\right)-1\right]$ (110)
It shows that the $q$ dispersion of $\tau_{SO}\left(q\right)$ suppresses the magnitude of $\sigma_{DOS}^{LV}$ to about $1/2$ of Eq.35
Appendix E The CPFs density
An expression for the CPFs density in terms of a generalized form of the fluctuation energy function $\Phi\left(q;H\right)$, which is valid at large wavenumbers, is given by (see Ref.MZJPC2023 ):
$n_{S}\left(H\right)=\left(\frac{7\zeta\left(3\right)E_{F}}{4\pi^{2}k_{B}T}\right)\frac{\pi}{d}\int_{0}^{q_{c}^{2}}\frac{d\left(q^{2}\right)}{\left(2\pi\right)^{2}}\frac{1}{\Phi\left(q;H\right)}$
where $\Phi\left(q;H\right)$ is given by Eq.39. Note the use of the general expression, Eq.39, for the dimensionless energy $\Phi\left(q;H\right)$ under the integral for $n_{S}\left(H\right)$, with the field-independent kinetic energy variable $x$, suggesting that the cutoff wavenumber $q_{c}$ should also be field independent.
Under the dirty limit condition: $\hbar/\tau_{SO}=\varepsilon_{SO}>>\mu_{B}H$, one finds:
$f_{+}\simeq\frac{\varepsilon_{SO}}{4\pi k_{B}T}\left[1+\left(1-2\left(\frac{\mu_{B}H}{\varepsilon_{SO}}\right)^{2}\right)\right]\simeq\frac{\varepsilon_{SO}}{2\pi k_{B}T},$
$f_{-}\simeq\frac{\varepsilon_{SO}}{4\pi k_{B}T}\left[1-\left(1-2\left(\frac{\mu_{B}H}{\varepsilon_{SO}}\right)^{2}\right)\right]\simeq$
$\frac{\varepsilon_{SO}}{2\pi k_{B}T}\left(\frac{\mu_{B}H}{\varepsilon_{SO}}\right)^{2}<<f_{+},$
and for $a_{\pm}$: $a_{+}\simeq 1,a_{-}\simeq\left(\frac{\mu_{B}H}{\varepsilon_{SO}}\right)^{2}<<1$, so that:
$\Phi\left(q;H\right)\simeq\Phi\left(x,\frac{T_{H}}{T}\right)\equiv\varepsilon_{H}+\psi\left(\frac{1}{2}+x+\frac{T_{H}}{T}\right)-\psi\left(\frac{1}{2}+\frac{T_{H}}{T}\right)$ (111)
where $T_{H}$ is given in Eq.75.
The linear approximation:
$\Phi\left(x,\frac{T_{H}}{T}\right)$ $\simeq$ $\Phi_{L}\left(x,\frac{T_{H}}{T}\right)=\varepsilon_{H}+\psi^{\prime}\left(\frac{1}{2}+\frac{T_{H}}{T}\right)x$ (112)
$\simeq\varepsilon_{H}+\eta\left(H\right)x$
is identical to the energy denominator of the fluctuation propagator in Eq.41.
The exact expression for the CPFs density is, then, written in the form:
$n_{S}\left(H\right)=n_{S}^{0}\int_{0}^{x_{c0}T_{c0}/T}\frac{dx}{\Phi\left(x,\frac{T_{H}}{T}\right)}$ (113)
where $n_{S0}$ is given in Eq.78.
Performing the integration in Eq.113 with $\Phi\left(x,T_{H}/T\right)$ replaced by the linear approximation $\Phi_{L}\left(x,T_{H}/T\right)$ we find for the linear approximation of the CPF density:
$n_{S}^{L}\left(H\right)=\frac{n_{S}^{0}}{\psi^{\prime}\left(\frac{1}{2}+\frac{T_{H}}{T}\right)}\ln\left(1+\frac{\widetilde{x}_{c}\left(H\right)}{\varepsilon_{H}}\right)$ (114)
where:
$\widetilde{x}_{c}\left(H\right)\equiv\psi^{\prime}\left(\frac{1}{2}+\frac{T_{H}}{T}\right)\left(\frac{T_{c0}}{T}\right)x_{c0}$
Using the asymptotic form of the digamma function for the low temperature limit $T/T_{H}\ll 1$ we find:
$n_{S}^{L}\left(H\right)$$\rightarrow$ $\left(\frac{T_{H}}{T}\right)n_{S}^{0}\ln\left(1+\frac{\widetilde{x}_{c}\left(H\right)}{\varepsilon_{H}}\right),$ (115)
$\widetilde{x}_{c}\left(H\right)$$\rightarrow$ $\left(\frac{T_{c0}}{T_{H}}\right)x_{c0}$
which is identical to Eq.77, provided $\widetilde{x}_{c}\left(H\right)$ is selected to be field independent, e.g.: $\widetilde{x}_{c}\left(H\right)\rightarrow 1$. The latter selection is consistent with the selection of the cutoff wavenumber to depend on field according to: $q_{c}=\xi^{-1}\left(H\right)$ (see Eq.69 and text around Eq.73).
We note that in our fitting procedure, employed in Refs.MZPRB2021 , MZJPC2023 , we have used the field-independent cutoff selection for $q_{c}$. Specifically, the best fitting value was found to agree with $x_{c0}=0.015$ (see Eq.87) with the selected values of the other parameters: $T_{c0}=0.212\boldsymbol{\mathit{K}}$ and $\varepsilon_{SO}=3\times 10^{-3}\boldsymbol{\mathit{eV}}$.
As indicated in the main text, deviations from the linear approximation may be significant for large cutoff wavenumbers which do not satisfy condition 86. Under these circumstances and in the low temperatures region, $T/T_{H}\ll 1$, we may use the asymptotic form of the digamma functions in Eq.111, which can be rewritten as: $\Phi\left(x,\frac{T_{H}}{T}\right)\rightarrow\varepsilon_{H}+\ln\left(x+T_{H}/T\right)-\ln\left(T_{H}/T\right)=\varepsilon_{H}+\ln\left(1+xT/T_{H}\right)$, that is:
$\Phi\left(x,\frac{T_{H}}{T}\right)$$\rightarrow$ $\varepsilon_{H}+\ln\left(1+\frac{\hbar Dq^{2}}{4\pi k_{B}T_{H}}\right),$ (116)
$T/T_{H}$ $\ll$ 11 11
In Fig.3 we plot $n_{S}\left(H\right)$ using Eq.113 with the exact expression, Eq.111, for $\Phi\left(x,T_{H}/T\right)$, together with $n_{S}^{L}\left(H\right)$ given in Eq.114, as functions of $T_{c0}/T$ for different values of the field $H$ and cutoff parameter $x_{c0}$.
Appendix F The TDGL AL conductivity
The AL conductivity is calculated within the TDGL functional approach from the time-ordered current-current correlator given in Eq.15, which can be rewritten in the form:
[…]
×$\times\sum\limits_{k=0,\pm 1,\pm 2,....}\frac{\Phi^{\prime}\left(x+\left|k+\nu\right|/2\right)}{\Phi\left(x+\left|k+\nu\right|/2\right)}\frac{\Phi^{\prime}\left(x+\left|k\right|/2\right)}{\Phi\left(x+\left|k\right|/2\right)}$
This is done by the analytic continuation from the imaginary Matsubara frequency $i\Omega_{\nu}$ to the real frequency $\omega$ , i.e.: $Q_{AL}\left(i\Omega_{\nu}\right)\rightarrow Q_{AL}^{R}\left(\omega\right)$, so that in the static limit: $\sigma_{AL}^{TDGL}=\lim_{\omega\rightarrow 0}\left(i/\omega\right)\left[Q_{AL}^{R}\left(\omega\right)-Q_{AL}^{R}\left(0\right)\right]$.
It is interesting to note that under direct analytic continuation of the discrete summation in Eq.117 about zero frequency, i.e. $\nu\rightarrow\hbar\omega/2\pi ik_{B}T\rightarrow 0$, all nonzero Matsubara-frequency terms are cancelled out and the remaining $k=0$ term can be written in the form:
[…]
Exploiting the linear approximation, i.e.:  $\Phi\left(x\right)\simeq\varepsilon_{H}+\eta\left(H\right)x$, and performing the integration over $x$ we find:
$\sigma_{AL}^{TDGL}\simeq\left(\frac{e^{2}}{16\hbar d}\right)\frac{\widetilde{\eta}\left(H\right)}{\varepsilon_{H}\left(1+\frac{\varepsilon_{H}}{\eta\left(H\right)x_{c}}\right)}$ (119)
in full agreement with Eq.49. Note the factor of $1/2$ multiplying the integer variables $\left|k+\nu\right|$ and $\left|k\right|$ in Eq.117, which was erroneously missing in a similar expression for this correlator in Refs.MZPRB2021 ,MZJPC2023 and consequently leading to an error by a factor of $2$ in the calculation of the AL conductivity there.
Appendix G CPFs localization length
Starting with Eq.91 for the correlation function $g\left(\rho/\xi_{0}\right)$ we will be interested here in its asymptotic behavior for $\rho/\xi_{0}\gg 1$, which will enable us to clearly identify the parameters that determine the localization of CPFs. Exploiting the linear approximation:
$\Phi\left(q;H\right)\simeq\widetilde{\varepsilon}_{H}+\frac{2}{\pi^{2}}\psi^{\prime}\left(1/2+\frac{T_{H}}{T}\right)\frac{T_{c0}}{T}\xi_{0}^{2}q^{2}$ (120)
and denoting: $\zeta\equiv\widetilde{q}/\sqrt{t},t\equiv T_{c0}/T$, we have:
[…]
Using the asymptotic form of the Bessel function, $J_{0}\left(z\right)\sim\left(2\pi z\right)^{-1/2}e^{\pm iz\mp i\pi/4}$ under the integral in Eq.121 and focusing only on the pole contributions at:
$\zeta=\pm i\zeta_{pole}=\pm i\sqrt{\frac{\pi^{2}}{2}\frac{\widetilde{\varepsilon}_{H}}{\psi^{\prime}\left(\frac{1}{2}+\frac{T_{H}}{T}\right)}}$
we may estimate the asymptotic behavior:
$g\left(\rho/\xi_{0}\right)\sim A\frac{\exp\left[-\left(\rho/\xi_{0}\right)\sqrt{\frac{\pi^{2}}{2}\frac{T}{T_{c0}}\frac{\widetilde{\varepsilon}_{H}}{\psi^{\prime}\left(\frac{1}{2}+\frac{T_{H}}{T}\right)}}\right]}{\left(\frac{T}{T_{c0}}\widetilde{\varepsilon}_{H}\right)^{1/4}\left[\psi^{\prime}\left(\frac{1}{2}+\frac{T_{H}}{T}\right)\right]^{3/4}}$ (122)
where $A$ is an adjustable parameter of the order one (see Fig.4).
It is dominated by the decaying exponential with the characteristic length:
$\rho_{loc}\left(H\right)=\xi_{0}/\sqrt{\frac{\pi^{2}}{2}\frac{T}{T_{c0}}\frac{\widetilde{\varepsilon}_{H}}{\psi^{\prime}\left(\frac{1}{2}+\frac{T_{H}}{T}\right)}}$
which is identical to Eq.92 in the main text. Its well-defined low-temperature limit can be easily found from the asymptotic form of the digamma function, $\psi^{\prime}\left(1/2+T_{H}/T\right)\sim T/T_{H}$ , that is:
$\rho_{loc}\left(H\right)$$\rightarrow$ $\left[\frac{2}{\pi^{2}}\left(\frac{T_{c0}}{T_{H}}\right)\frac{1}{\widetilde{\varepsilon}_{H}}\right]^{1/2}\xi_{0};$
$T$ $\ll$ $T_{H}$